### v2

# Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments. The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”

Abstract
I study the optimal allocation of positional goods in the presence of externalities arising from consumers’ concerns for relative consumption. Using a mechanism design approach, I characterize the externalities by a feasibility condition. I show that the revenue-maximizing mechanism possibly excludes some buyers and fully separates the participants if their type distribution satisfies Myerson’s regularity. The seller can guarantee at least half the maximum revenue by offering a single level of positional goods. When there is no exclusion, as the seller offers more positional good levels, the consumer surplus decreases (increases) if the distribution has an increasing (decreasing) failure rate. Higher participation increases the consumer surplus under increasing failure rates. Applications include luxury goods, priority services, education, and organizational hierarchies.
Keywords: Positional goods, mechanism design, externalities, feasibility.
If everyone stands on tiptoe, no one sees better.
— Fred Hirsch (1976), Social Limits to Growth
1 Introduction
Positional goods are goods or services whose value to consumers depends on their relative positions in consumption. Hirsch (1976) first coined the concept of positional good that is “either scarce in some absolute or socially imposed sense” or “subject to congestion or crowding through more extensive use” (pp. 27). Because consumers care about their relative positions, allocating an item to one buyer has externalities on others.
Examples of positional goods abound. Consumer goods, especially luxury goods, are often positional because people derive happiness from comparisons (Frank, 1985a; Luttmer, 2005; Carlsson et al., 2007). Another example is priority services and front-running (Gershkov and Winter, 2023), such as priority boarding. Such services reduce consumers’ waiting time by moving them ahead of the crowd, but they become less effective as more consumers use them. Additionally, education is the classic example in Hirsch (1976): “The value to me of my education depends not only on how much I have but also on how much the man ahead of me in the job line has.”
More broadly, status within an organization can also be viewed as a positional good, as employees invest efforts to move up in the hierarchy (Moldovanu, Sela, and Shi, 2007). As a revenue-maximizing seller, the organization designer can also induce effort through employees’ positional concerns and exclude low-ability employees.
In these examples, the allocation of positional goods has externalities on consumers, as moving up the position of one consumer would push down that of another. At the same time, when more consumers buy the same good (e.g., priority boarding), its value depreciates due to more extensive use. Thus, unlike ordinary goods, it is impossible to allocate the good (e.g., the first place in line) to multiple consumers at face value: “What each of us can achieve, all cannot” (Hirsch, 1976). In general, positional goods induce a game of status, in which agents pay in money or effort to attain higher relative positions. 111 Because of the dual interpretation of the model, I use “consumers” and “agents”, as well as “seller” and “principal” interchangeably. While the principal can exploit the status game to incentivize spending or effort, this often comes at the expense of agent welfare. 222 For example, Asian governments have been orchestrating educational disarmament (The Economist, 2021). U.S. colleges switched to coarse grading, such as pass/fail, to relieve student stress during the COVID-19 pandemic. Luxury consumption is discouraged by many cultures and religions.
In this paper, I study the optimal allocation of positional goods using a mechanism design approach. I assume buyers care about their relative positions, or status, which depends on the mass of consumers who buy goods at the same level and at lower levels. Moreover, they have private information about their types—i.e., valuations of the status (or marginal costs of effort). The seller can offer multiple levels of positional goods at different prices. The buyer can pay a price (or effort) for one of the goods to obtain the status value, or pay nothing and opt out.
First, I characterize positional externalities by a feasibility condition: the interim status allocation is feasible (i.e., can be induced by an allocation of positional good) if and only if it is weakly majorized by the buyer’s type distribution in the quantile space.
Then, I focus on revenue maximization. Using results in Kleiner, Moldovanu, and Strack (2021, henceforth KMS), I show that offering more levels of positional goods increases the seller’s revenue if the type distribution is regular in the sense of Myerson (1981). Therefore, the revenue-maximizing mechanism excludes types with negative marginal revenue and fully separates participants, which can be implemented by an all-pay auction with a reserve price, where the buyer who pays more obtains the higher position. In contrast, if Myerson’s regularity condition fails for some types, the optimal mechanism pools them into the same level of positional goods.
I also find that a simple posted-price mechanism that offers a single level of positional goods can guarantee at least half the maximal revenue. In the proof, I consider an auxiliary screening problem of selling a continuum of ordinary (i.e., nonpositional) goods to a continuum of buyers. The screening problem highlights the distinction between selling positional and ordinary goods: in the former, allocating the good to one buyer has externalities on others, so the seller cannot sell the highest status to multiple buyers at face value. Therefore, selling positional goods results in lower revenue compared to selling ordinary goods (Mussa and Rosen, 1978).
Next, I study how the number of positional good levels and exclusion affect consumer welfare. I first focus on the case when consumer participation is fixed, for example, when there is a free regular service available. I show that the consumer surplus is higher when the seller offers more positional good levels if the buyer’s type distribution has an increasing failure rate (IFR), which implies a thin tail. This result highlights the tension between the seller’s revenue and the consumer surplus. On the other hand, when the type distribution has a decreasing failure rate (DFR), which implies a heavy tail, the consumer surplus is higher when the seller offers fewer positional good levels. Furthermore, under IFR, expanding consumer participation increases the consumer surplus, as the marginal gain of additional participants outweighs the inframarginal loss of the existing participants from reduced status due to pooling.
I also study optimal mechanisms that maximize the consumer surplus. When negative transfers are allowed, full separation without exclusion maximizes the consumer surplus. Intuitively, this is because assortative matching, and possibly negative transfers can redistribute the efficiency gains among consumers (see also Gershkov and Schweinzer, 2010). By contrast, in applications where the price (effort) must be nonnegative, such as organizational hierarchies and education, I show that total pooling without exclusion maximizes the consumer surplus under IFR. Unlike selling ordinary goods, exclusion raises the status of participants who would be otherwise pooled with the excluded ones. Nevertheless, under IFR, the harm of exclusion to the excluded consumers exceeds its benefit to the participants. In contrast, if the distribution has a decreasing failure rate or a heavy tail (e.g., the Pareto distribution), the consumer surplus under the nonnegative price constraint is maximized by full separation without exclusion.
For social welfare, if the social planner puts the same weight on the seller’s revenue and the consumer surplus, then full separation without exclusion is optimal because it leads to assortative matching, and transfer is welfare-neutral.
As an extension, I consider the case where the intrinsic value of the positional good is determined by its intrinsic quality produced at a cost (à la Mussa and Rosen (1978)), and the results remain robust. Moreover, I also allow the seller to deliberately lower the status, or equivalently, prolong the waiting time by pausing service or procrastinating. Thus, the seller can create negative status or excessive waiting time as a substitute for exclusion. I show that the revenue-maximizing mechanism fully separates those with positive virtual valuations and assigns negative status to those who with negative virtual valuations. Additionally, I consider cases where the status game is a “suffering” in that higher types suffer more from participation, and the results remain robust.
The model has dual interpretations that lead to two strands of applications. On the one hand, it is a monopolist problem of selling positional goods, such as priority services (Gershkov and Winter, 2023) and luxury goods. Alternatively, it can be interpreted as an organizational design problem where the principal motivates agent effort (e.g., Moldovanu, Sela, and Shi, 2007; Immorlica, Stoddard, and Syrgkanis, 2015) or maximizes agent welfare through status within an organization. Under this interpretation, the optimal mechanism that maximizes effort possibly excludes some types and fully separates the participants if the ability distribution satisfies Myerson’s regularity. This can be seen in “rank-and-yank” organizations that rank every employee and terminate those who underperform.
The implication also echoes the education example by Hirsch (1976) and sheds light on the consequences of education arms races. From the agent’s perspective, while the common wisdom calls for educational disarmament and less intense competition, I show that this benefits agents if their ability distribution has a thin tail. Under the IFR property, admission lotteries or random assignments, instead of meritocratic allocation, maximizes agent welfare. On the contrary, if the ability distribution has a heavy tail, meritocracy maximizes agents’ welfare (and their effort). Additionally, if education is sufficiently socially productive, educational arms races can be optimal for social welfare.
In Appendix A, I consider the case where exclusion is impossible, which is true in many applications where the lowest status is free. For example, a society may not allow exclusion, so investing no effort guarantees the lowest social status (Moldovanu, Sela, and Shi, 2007). Moreover, priority services may be offered in addition to free regular services (Gershkov and Winter, 2023). In these cases, the lowest-level positional good may still have a positive value, as the lowest status cannot be offered to multiple agents at face value either. I provide the necessary and sufficient conditions for the revenue-maximizing mechanisms to be separating or pooling and for consumer welfare to be decreasing or increasing in the number of positional good levels, as well as some additional results, such as approximation by offering two levels. The results are qualitatively similar to the benchmark model with exclusion.
Literature Review.
The idea that the consumer’s utility depends on the comparison with others’ consumption dates back to Veblen (1899) (see also Duesenberry, 1949; Leibenstein, 1950; Bagwell and Bernheim, 1996). Since Hirsch (1976) introduced the concept of positional good, the status interdependence in preference has been formally modeled and studied by Frank (1985b, 2005, 2008), Robson (1992), and Hopkins and Kornienko (2004).333 Another strand of literature uses the Duesenberry formulation that stems from Duesenberry (1949) and Pollak (1976) (see the survey by Truyts (2010)), which assumes the consumer’s utility depends on the consumption of others in addition to their absolute consumption. This literature focuses on consumer choice in the presence of positional externalities and the welfare effects of changes in the income distribution and government policies (e.g., luxury tax).
Empirical evidence of positional goods and externalities include Frank (1985a, 2005), Luttmer (2005), Alpizar et al. (2005), Carlsson et al. (2007), Bursztyn et al. (2018). In particular, education is a typical example of positional good (Hirsch, 1976). Durst (2021) finds empirical evidence that education is indeed positional in Germany.
Moldovanu, Sela, and Shi (2007, henceforth MSS) study the optimal design of the number and size of status categories in an organization to maximize agents’ effort when they care about their relative position (i.e., status). They find that full separation is optimal if the ability distribution satisfies IFR, while coarse partition may be optimal depending on the distribution—for example, the optimal partition involves only two categories if it is sufficiently concave. Despite the similarities in our results, this paper is different in three ways.444 The definition of status in MSS, which equals the number of strictly inferior agents minus the number of strictly superior agents, is equivalent (up to an affine transformation) to my specification (see Footnote 10). First, I allow for exclusion of agents and stochastic allocations and study agent welfare. In reality, the exclusion of agents is common in organizations, and employee welfare is an important consideration. Second, in the case without exclusion, I obtain stronger results that are either necessary and sufficient conditions or require fewer distributional assumptions (see Appendix A). Third, I use a mechanism design approach with a feasibility characterization using majorization, while they use methods in optimal contest design (with finite agents).
Relatedly, Rayo (2013) studies a model of monopolist provision of conspicuous goods and characterizes the optimal allocation. His results are qualitatively similar to MSS’s, and he does not consider exclusion, consumer welfare maximization, or stochastic mechanisms either. Conspicuous goods in his definition also differ from positional goods (or status in MSS) in that consumers care about the average type in their assigned category instead of relative positions. Nevertheless, the feasibility condition can also be applied to conspicuous goods in his model (see Remark 4), which explains the similarities between his results and those of positional goods.555 Board (2009) assumes more general benefits from being in a category (i.e., group quality) and finds a contradictory result that the revenue-maximizing principal segregates agents too finely compared to the efficient level (which is full separation in my model). He also finds that revenue maximization leads to too much exclusion.
Similar to MSS, Immorlica, Stoddard, and Syrgkanis (2015) study the optimal design of badges to incentivize users’ contribution (i.e., effort) to the online community through the status value. Unlike MSS, they consider exclusion and allow the status value to be concave or convex in the relative position. They find the “leaderboard with a cutoff” mechanism is optimal if the distribution satisfies Myerson’s regularity. As in MSS, they do not consider agent welfare or stochastic allocations, which is an important consideration in online communities. Most importantly, their status value function is strictly decreasing in the number of people weakly superior to them (without adjustment to those at the same level), which naturally precludes pooling from being optimal because it decreases the status of every agent pooled together. Consequently, (i) total pooling will never be optimal—even for agent welfare—in sharp contrast to my result, (ii) the optimality of full separation in effort maximization is less surprising, and (iii) exclusion is equivalent to pooling at the lowest status.
In a different vein, Gershkov and Winter (2023, henceforth GW) study a model of priority service and analyze its implications for consumer welfare. Priority service (PS) is an example of positional goods that arise from technological constraints. Using a mechanism design approach, this paper generalizes their model by allowing for multiple (and stochastic) PS levels and exclusion. By contrast, they focus on two PS levels instead of the optimal mechanism. 666 They extend their model to multiple PS levels but without exclusion in Section V. In the case without exclusion, I refine some of their results by providing necessary and sufficient conditions and show some additional results in Appendix A.
Another closely related paper is by Loertscher and Muir (2022), who study revenue maximization when selling a fixed inventory of goods with heterogeneous qualities. Without the intrinsic value and under linear status, my model is equivalent to a continuous version of theirs where each quality has a unit stock. In particular, pooling multiple types into the same status corresponds to randomization (or conflation) in their model.
More broadly, this paper is also related to mechanism design with allocative externalities (e.g. Jehiel, Moldovanu, and Stacchetti, 1996; Jehiel and Moldovanu, 2005; Dworczak, Reuter, Kominers, and Lee, 2024).
2 Model
2.1 Setup
A monopolist seller (she) sells positional goods $q\in X\subseteq\mathbb{R}_{++}$ to a continuum of buyers (he) with unit mass. Denote by $q=0$ the outside option of not buying the good. Because consumers care about their relative position, consumption of positional goods induces buyers’ status (as defined later). Buyers are heterogeneous in their valuations of status (i.e., types), denoted by $\theta$. Assume the buyer’s type $\theta$ has a distribution $F(\theta)$ with full support $\Theta=[0,\bar{\theta}]$ and density $f(\theta)$.777 The support can be extended to $\bar{\theta}=\infty$ and $\underline{\theta}>0$. The latter case leads to a complication in consumer surplus maximization under decreasing failure rates (see Remark 11). A buyer with type $\theta$ who pays $p$ for a status $s\in[0,1]$ from the good receives a payoff of $u(p,s,\theta)=\theta s-p+v(\theta)$, where $v(\theta)\geq 0$ is the intrinsic value of the good. 888 I assume for now that the intrinsic value is independent of $q$ to abstract from quality differentiation (Mussa and Rosen, 1978). In Section 4.1, I consider an extension where the positional good also has an intrinsic quality equal to ${q}$ produced at a cost $c({q})$. Assume $v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$, so higher types have higher intrinsic values. 999 In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$, so higher types have lower intrinsic values. If the buyer does not get the good, his status is defined as $s=0$, and he does not get the intrinsic value.
Definition of status.
Denote by $G$ the distribution of buyers’ consumption of the good $q$ over $\tilde{X}\equiv X\cup\{0\}$. Then, $G({q})$ represents the mass of buyers who consume weakly lower-level goods than ${q}$ (e.g., smaller cars, lower boarding groups) or opt out ($q=0$). Following Robson (1992) and Hopkins and Kornienko (2004), I define the status of a buyer who consume $q\in X$ as
$S({q},G(\cdot))=\frac{G^{-}({q})+G({q})}{2}\in[0,1],$ (1)
where $G^{-}({q})\equiv\lim_{{q}\to{q}^{-}}G({q})$ is the mass of buyers who consume strictly lower-level goods than ${q}$ (or opt out). In other words, the status equals the mass of buyers who consume strictly lower-level goods plus half the mass of buyers at the same level.101010 Hopkins and Kornienko (2004) assume a more general specification: $S(q,G(\cdot))=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1)$. If $\gamma>1/2$ ($\gamma<1/2$), pooling will decrease (increase) the total expected status. I will show that many results are robust to the general specification. An equivalent specification (up to an affine transformation), à la MSS and Dubey and Geanakoplos (2010), is $S({q},G(\cdot))=G^{-}({q})-(1-G(q))\in[-1,1],$ which results in an expected total status of 0. Buyers who opt out ($q=0$) are considered as strictly below those who buy $q\in X$.
From another perspective, because buyers only care about the induced status, the seller essentially sells status $s\in[0,1]$, with distribution $H$, subject to the feasibility constraint
$s=\frac{H^{-}({s})+H({s})}{2}\in[0,1],\text{ for all }{s}\in\operatorname{supp}(H)\setminus\{0\}.$ (2)
Those who do not buy get status $s=0$. The expected total status is always a constant (equal to half the mass of participants) regardless of the mass points of the distribution, corresponding to the zero-sum nature of the status game.
This specification of status can arise from interpersonal comparisons (Frank, 1985b; Robson, 1992; Hopkins and Kornienko, 2004), as buyers derive utility from consumers of lower-level goods (e.g., smaller cars), but not as much from those of the same level. Moreover, it can also arise from technological constraints. For example, the agent’s education level determines his place in the job line (or social status), and his payoff depends on the number of people ahead of him and at the same place in the line (or the hierarchy). Below is another example in queuing.
Example 2.1 (Priority Services (Gershkov and Winter, 2023)).
A unit mass of customers arrive at the same time. A seller can serve one customer at each instant and offers priority services to skip the line: consumers at higher priority levels are served first, and those of the same level are served in random order. The consumer’s payoff is $u(p,t,\theta)=\tilde{v}(\theta)-\theta t-p$, where $t$ is the waiting time, $\theta$ is the value of time (or the cost of waiting), and $\tilde{v}(\theta)$ is the value of the service. This is equivalent to the model with $s=1-t$ and $v(\theta)=\tilde{v}(\theta)-\theta$. 111111 The assumption $v^{\prime}(\theta)\geq 0$ requires $\tilde{v}^{\prime}(\theta)\geq 1$ here, that is, higher types derive higher utility from the service after accounting for waiting costs. In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$ (i.e., $\tilde{v}^{\prime}(\theta)\leq 0$) so that higher types derive higher utility from the service after accounting for waiting costs. The seller essentially sells the waiting time $t\in[0,1]$ subject to the feasibility constraint
$t=\frac{\Phi^{-}({t})+\Phi({t})}{2}\in[0,1],$
where $\Phi$ is the distribution of the waiting time $t$.
Because the status category within organization is also a positional good, another example is the status contests in organizations, where the principal (seller) maximizes effort by designing status categories of agents (buyers).
Example 2.2 (Status Contests (Moldovanu et al., 2007)).
Consider a principal who maximizes agent effort by designing status categories. Agents care about their relative positions within the organization. Then, the positional good $q$ represents the status category, the price $p\geq 0$ represents the agent’s effort, and the type $\theta$ represents the agent’s ability (i.e., marginal cost of effort). The agent has a linear effort cost $p/\theta$ and a payoff $s-p/\theta$ if he has status $s\in[0,1]$ in the organization. When the agent is excluded or expelled, he receives zero payoff. This is a special case of the model with $v(\theta)=0$ and $p\geq 0$ after scaling the payoff by $\theta$.
2.2 A Mechanism Design Approach
Consider a direct mechanism $(p(\theta),\chi(\theta,\omega))$, consisting of a payment function $p\colon\Theta\to\mathbb{R}$ and a (potentially stochastic) positional good allocation $\chi\colon\Theta\times\Omega\to\tilde{X}=X\cup\{0\}$, where the random variable $\omega\in\Omega=[0,1]$ captures the randomness in the mechanism.121212 The allocation can be implemented by a pricing scheme as a function of the distribution of positional good levels. An allocation $\chi$, together with the type distribution $F$, induces a distribution of positional goods $q$, given by $G_{\chi}(q\mid\omega)=\int\mathbf{1}[\chi(\theta,\omega)\leq q]\mathop{}\!\mathrm{d}F(\theta)$.
Thus, it is equivalent to consider a direct mechanism $\{p(\theta),s(\theta)\}$, where $s\colon\Theta\to[0,1]$ is the interim status induced by the allocation $\chi$, given by
$s(\theta)=\operatorname{\mathbf{E}}_{\omega}[S(\chi(\theta,\omega),G_{\chi}(\cdot\mid\omega))].$
In particular, under a deterministic allocation $\chi\colon\Theta\to\tilde{X}$, the interim status is $s(\theta)=S(\chi(\theta),G_{\chi})$.
Denote by $U(\theta)=\theta s-p(\theta)+v(\theta)$ the buyer’s payoff if he participates and $U(\theta)=0$ otherwise. 131313 Although stochastic mechanisms are allowed, it is never optimal to randomly exclude participants either for the seller or consumers. This can be formally shown by multiplying $U(\theta)$ (or $v(\theta)$) by an expected participation dummy $\delta(\theta)\in[0,1]$and showing that $\delta^{*}(\theta)=1$ for all $\theta\geq\theta_{0}$. The following lemmas are implied by standard arguments.
Lemma 1.
There exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates (i.e., $U(\theta)\geq 0$) if and only if $\theta\geq\theta_{0}$.
Lemma 2.
A direct mechanism $\{p(\theta),s(\theta)\}$ is incentive-compatible if and only if
  • $U(\theta)=U(\theta_{0})+\int_{\theta_{0}}^{\theta}(s(x)+v^{\prime}(x))\mathop{}\!\mathrm{d}x$ for all $\theta\in[\theta_{0},\bar{\theta}]$, and
  • $s(\theta)$ is increasing.
An incentive-compatible interim status $s(\theta)$ is not necessarily feasible, as it may be unable to be induced by an allocation $\chi$. Formally, say an interim status $s(\theta)$ is feasible if there exists an allocation $\chi$ that induces it—i.e., $s(\theta)=\operatorname{\mathbf{E}}_{\omega}[S(\chi(\theta,\omega),G_{\chi}(\cdot\mid\omega))]$.
To provide intuitions, consider a deterministic allocation without exclusion $\chi\colon\Theta\to{X}$. Incentive compatibility implies that $\chi(\theta)$ is increasing because $s(\theta)=S(\chi(\theta),G_{\chi})$ is increasing. When $\chi(\theta)$ is strictly increasing at $\theta$, then $s(\theta)=F(\theta)$. Otherwise, when $\chi(\theta)=q>0$ is constant on some interval $[\theta_{1},\theta_{2}]$ (and $\chi(\theta)\neq q$ outside this interval), then
$s(\theta)=\frac{G^{-}(q)+G(q)}{2}=\frac{F(\theta_{1})+F(\theta_{2})}{2}=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\tilde{\theta}\in[\theta_{1},\theta_{2}]]=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\chi(\tilde{\theta})=\chi(\theta)]$
for all $\theta\in[\theta_{1},\theta_{2}]$. In either case, $s(\theta)=S(\chi(\theta),G_{\chi})=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\chi(\tilde{\theta})=\chi(\theta)]$.
Now I characterize the necessary and sufficient condition for an increasing interim status $s(\theta)$ to be feasible in general.
Theorem 1 (Feasibility).
An increasing $s$ is feasible if and only if it is weakly majorized by $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}_{w}(F)$, that is,
$\int_{x}^{\bar{\theta}}s(\theta)\mathop{}\!\mathrm{d}F(\theta)\leq\int_{x}^{\bar{\theta}}F(\theta)\mathop{}\!\mathrm{d}F(\theta)\text{ for all }x\in[0,\bar{\theta}].$ (3)
The proof is à la Theorem 3 (Border’s Theorem) in KMS.
Proof.
($\implies$) follows from $\operatorname{\mathbf{E}}[s(\theta)\mid\theta\geq x]\leq\operatorname{\mathbf{E}}[F(\theta)\mid\theta\geq x]$ for all $x\in[0,\bar{\theta}]$ because switching from any $s(\theta)$ to assortative rule ($s(\theta)=F(\theta)$) takes the status of low type to high types.
[…]
$s(\theta)=\frac{1}{F(\bar{\theta}_{i})-F(\underline{\theta}_{i})}\int_{\underline{\theta}_{i}}^{\bar{\theta}_{i}}F(\theta)\mathop{}\!\mathrm{d}F(\theta)=\frac{F(\underline{\theta}_{i})+F(\bar{\theta}_{i})}{2};$
and (iii) fully separating otherwise, where $s(\theta)=F(\theta)$.
Hence, by Proposition 1 (Choquet’s Theorem) in KMS, measure $\lambda_{s}$ supported on the extreme points of $\operatorname{\text{MPS}}_{w}(F)$ such that $s=\operatorname{\mathbf{E}}[\tilde{s}\mid\tilde{s}\sim\lambda_{s}]$. Thus, $s$ can be induced by randomizing over allocations that each induces extreme points of $\operatorname{\text{MPS}}_{w}(F)$. ∎
Remark 1.
For deterministic allocations, $s(\theta)$ must be an extreme point of $\operatorname{\text{MPS}}_{w}(F)$—i.e., either the majorization constraint or the monotonicity constraint binds on $[\theta_{0},\theta]$ and $s(\theta)=0$ for all $[0,\theta_{0}]$ (see Theorem 1 and Corollary 2 in KMS). For stochastic allocations, $s(\theta)$ can take other forms.
Remark 2.
When exclusion is impossible, the equality must hold at $x=0$, that is, $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[F]=1/2$. In other words, $s(\theta)$ is a mean-preserving spread of $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}(F)$.141414 Without exclusion, one can also write $s(\theta)=\operatorname{\mathbf{E}}_{{q}\sim\chi(\theta,\Omega)}[\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid{q}]]$, where $\chi(\theta,\Omega)$ denotes the distribution of $q=\chi(\theta,\omega)$ given $\theta$ generated by the randomness $\omega\in\Omega$.
Remark 3.
Under MSS’s specification $S({q},G)=G^{-}({q})-(1-G({q}))$ without exclusion, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(2F-1)$ with $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[2F-1]=0$.
Remark 4.
In Rayo’s (2013) model of conspicuous goods, where the status value induced by $\chi$ is $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$, the feasibility condition is $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$.
Intuitively, we have $s(\theta)=F(\theta)$ under full separation. In contrast, any (partial) pooling of multiple types into the same positional good makes the status $s(\theta)$ a mean-preserving spread of $F(\theta)$ in the quantile space because the value of the positional good becomes lower than its face value, or the value under full separation.
Example (Two Status Levels).
Suppose there are two levels of positional goods, low-level $(p_{L},s_{L})$ and high-level $(p_{H},s_{H})$, where $p_{L}=0$ (so there is no exclusion). Denote by $\theta^{*}$ the cutoff type whose is indifferent between the two (determined by $p_{H}$). The consumer buys $s_{H}$ if and only if $\theta\geq\theta^{*}$. Then, the interim status is
$s(\theta)=\begin{cases}s_{L}=F(\theta^{*})/2,&\text{ if }\theta<\theta^{*},\\ s_{H}=(1+F(\theta^{*}))/2,&\text{ otherwise}.\end{cases}$
It is straightforward to check that $\operatorname{\mathbf{E}}[s]=1/2$ and that $\int_{\theta}^{\bar{\theta}}s(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})\leq\int_{\theta}^{\bar{\theta}}F(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})$ for all $\theta\in[0,\bar{\theta}]$ (with equality at 00, $\theta^{*}$, and $\bar{\theta}$).
Figure 1: Two Priority Levels
3 Optimal Mechanisms
In this section, I consider the optimal mechanisms that maximize the seller’s revenue and study its welfare implications. Two extreme mechanisms are particularly of interest: full separation (i.e., $s(\theta)=F(\theta)$) and total pooling among participants (i.e., $s(\theta)=(1+F(\theta_{0}))/2$). The former can be induced by offering a continuum of positional goods that lead to a proliferation of status (i.e., the finest partition), and the latter can be induced by offering only one level of positional good.
3.1 Revenue Maximization
The revenue (or effort) maximization problem is given by
$\max_{s(\theta),p(\theta),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}p(\theta)\mathop{}\!\mathrm{d}F(\theta)$ (4)
subject to the following constraints for all $\theta\in[\theta_{0},\bar{\theta}]$:
$U(\theta)\equiv\theta s(\theta)-p(\theta)+v(\theta)\geq 0$ (IR) (5)
$U(\theta)=U(\theta_{0})+\int_{\theta_{0}}^{\theta}(s(x)+v^{\prime}(x))\mathop{}\!\mathrm{d}x$ (IC) (6)
$s(\theta)\mbox{ is increasing}$ (7)
$s\in\operatorname{\text{MPS}}_{w}(F)$ (MPS) (8)
Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. Say $F$ satisfies the Myerson’s regularity if $J(\theta)$ is increasing. Say $F$ has an increasing failure rate (IFR) if $\frac{1-F(\theta)}{f(\theta)}$ is decreasing. The revenue is given by
$R=\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)-U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$ (9)
[…]
$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (10)
The maximum is attained at the extreme point of $\operatorname{\text{MPS}}_{w}(F)$, which is an extreme point of $\operatorname{\text{MPS}}(F\cdot\mathbf{1}_{[\theta_{0},\bar{\theta}]})$. The following proposition follows from Fan-Lorentz Theorem (see, e.g., Theorem 4 in KMS).
Proposition 2.
If $J(\theta)$ is increasing, then
  1. (i)
    The revenue is increasing in the number of positional good levels.
  2. (ii)
    The revenue-maximizing mechanism excludes $\theta<\theta_{0}^{*}$ and fully separates $\theta\geq\theta_{0}^{*}$, where the optimal cutoff is $\theta^{*}_{0}=\operatorname*{arg\,max}_{\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$.
If $J(\theta)$ is decreasing on some $\Theta_{p}\subseteq[0,\bar{\theta}]$, then the revenue-maximizing mechanism pools every $\theta\in\Theta_{p}$ into the same positional good level.
The proof is in Appendix B.
Remark 5.
Proposition 2 still holds under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1/2]$. To see this, when $\gamma<1/2$ (e.g., Immorlica et al. (2015) where $\gamma=0$), pooling is even less likely to be optimal as it reduces expected total status. 161616 Formally, fix a positional good allocation $\chi$, the induced interim status $\tilde{s}(\theta)$ when $\gamma<1/2$ and $s(\theta)$ under the original specification ($\gamma=1/2$) satisfy $\tilde{s}(\theta)\leq s(\theta)$ for all $\theta\in[\theta_{0},\bar{\theta}]$, with equality if and only if $\tilde{s}(x)=F(x)$ for all $x\in[\theta_{0},\theta]$.
Remark 6.
Under the signaling specification $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$ (see Remark 4), Proposition 2 also holds (with a different optimal cutoff $\theta^{*}_{0}$) because the only difference is that $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$ instead.
Remark 7.
When restricted to deterministic mechanisms and under the specification $\tilde{S}({q},G)=\phi(S({q},G))$, where $\phi$ is an increasing and convex function, Proposition 2 still holds (with a different $\theta^{*}_{0}$) because Fan-Lorentz Theorem still applies (see details in the proof).
The revenue-maximizing mechanism can be implemented by an all-pay auction with a reserve price, where the higher the buyer pays (above the reserve price), the higher-level positional good he receives. In applications such as education or social status, it can be interpreted as an all-pay contest in which those whose effort does not meet the reserve price are punished with exclusion (expulsion from the organization).
If the virtual valuation $J(\theta)$ is decreasing for some types, the revenue-maximizing mechanism pools these types into the same positional good level. In general, the revenue-maximizing mechanism is derived by applying the ironing technique to obtain the convex hull of $\tilde{J}(\theta)\equiv\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$.
The proposition has several implications. First, under Myerson’s regularity condition, it is optimal to sell as many levels of positional goods as possible. This is because as the number of position levels increases, the seller can induce a finer partition of buyers and thus extract a higher revenue. 171717 For any two increasing allocations $s,\hat{s}\in\operatorname{\text{MPS}}(F)$ on $[\theta_{0},\bar{\theta}]$, if $s$ is finer than $\hat{s}$, then $\hat{s}\in\operatorname{\text{MPS}}(s)$ (the converse is not true). Thus, I also say $s$ is finer than $\hat{s}$ in terms of majorization if $\hat{s}\in\operatorname{\text{MPS}}(s)$. For example, airlines should introduce more boarding groups (or flight classes) to the extent regulations and practical constraints allow. Luxury companies should offer a proliferation of products from factory stores to high-end.
Second, compared to selling ordinary goods with intrinsic value $v(\theta)$ only, attaching the status value benefits the seller. This explains why luxury companies invest heavily in marketing to create the status value.
Third, the seller also benefits from excluding low types. For uniform distribution on $[0,1]$ and when $v(\theta)=0$, exclusion increases the revenue by 23.5%. Thus, exclusion is important for the seller (organization designer) to extract more revenue (effort). In applications where exclusion is impossible, full separation remains optimal under the regularity condition (See Proposition A.1). The following corollary focuses on the properties of the optimal exclusion level $\theta^{*}_{0}$.
Corollary 2.1.
Assume $F$ satisfies IFR. Then, we have the following:
  1. (i)
    The optimal exclusion $\theta^{*}_{0}\leq J^{-1}(0)$;
  2. (ii)
    If $v(0)\geq(v^{\prime}(0)+1)/f(0)$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=0$).
Remark 8.
If $v(\theta)=\alpha\theta$ for some $\alpha\geq 0$, then the optimal cutoff is $\theta_{0}^{*}=J^{-1}(0)$.
Because of the intrinsic value $v(\theta)$, under the IFR property, the revenue-maximizing mechanism here leads to lower exclusion than the standard result that excludes agents with negative virtual valuation $J(\theta)<0$. Furthermore, when the intrinsic value $v(0)$ is large enough and there are many low types ($f(0)$ is large), the seller cannot benefit from exclusion.
Approximation.
In some applications, the number of positional good levels the seller can offer is subject to regulation, practical constraints, or other costs. The following proposition shows that the seller can still obtain a substantial portion of the maximum revenue by offering only one level of positional good.
Proposition 3.
The seller can obtain at least half the maximum revenue by selling one level of positional good.
{proofsketch}
Consider an auxiliary screening problem of selling indivisible ordinary goods (e.g., apples), where a posted price $p^{*}=\operatorname*{arg\,max}_{p}p(1-F(p))$ is optimal. I show that (i) by selling one level of positional good, the seller can obtain at least $\frac{1}{2}\max p(1-F(p))$ because those who buy the good get at least a status level of $s=1/2$; and (ii) the maximum revenue from selling positional goods is strictly below $\max p(1-F(p))$ because of positional externalities. (See the complete proof in Appendix B.)
Remark 9.
The lower bound does not require IFR or Myerson’s regularity.
The auxiliary screening problem considered in the proof, which is equivalent to selling a continuum of indivisible ordinary goods to a continuum of buyers, highlights the distinction between selling ordinary goods and positional goods. In the screening problem, selling an item to one consumer has no externalities on others. By contrast, negative externalities are present when selling positional goods—the seller cannot offer the highest status to multiple buyers at face value. Thus, the seller of positional goods is subject to an additional constraint, resulting in strictly less revenue.
Graphical Illustration.
Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$. Figure 2 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$. Assume the intrinsic value $v(\theta)=0$ for simplicity.181818 A positive intrinsic value $v(\theta)\geq 0$ will increase $\frac{R_{1}}{\max R}$ by increasing both $R_{1}$ and $\max R$. In Figure 2, (i) the blue area in the left panel represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. (ii) The red area in the right panel represents the revenue from selling one level of positional good at price $F(\theta_{0}^{*})$—i.e., $\bar{s}(\theta)=\frac{1+F(\theta_{0}^{*})}{2}\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. 191919 To see (i) and (ii), the revenue under $s^{*}$ is $\int_{\theta_{0}^{*}}^{1}J(x)F(x)\mathop{}\!\mathrm{d}F(x)=-\int_{\tau^{*}_{0}}^{1}\tau\mathop{}\!\mathrm{d}R(\tau)=\tau^{*}_{0}R(\tau^{*}_{0})+\int_{\tau^{*}_{0}}^{1}R(\tau)\mathop{}\!\mathrm{d}\tau$. Analogously, the revenue under $\bar{s}$ is $\int_{\theta_{0}^{*}}^{1}J(x)\frac{1+F(\theta_{0}^{*})}{2}\mathop{}\!\mathrm{d}F(x)=\frac{1+\tau^{*}_{0}}{2}R(\tau^{*}_{0})$. Because the seller can optimize the price (and setting it at $F(\theta_{0}^{*})$ is feasible), the revenue from selling one level of positional good is at least the red area. (iii) The entire box in either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem. It is straightforward to see that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
Figure 2: Revenue curve $R(\tau)$ for the uniform distribution
The approximation is very close for many common distributions, as shown in the following examples (with $v(\theta)=0$).
Example (Exponential distribution).
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, selling one level of positional good can obtain $91.9\%$ of the maximum revenue.
Example (Uniform distribution).
If $F(\theta)=\theta$ on $[0,1]$, selling one level of positional good can obtain $92.4\%$ of the maximum revenue.
Example (Power distribution).
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, selling one level of positional good can obtain
$\frac{R_{1}}{\max R}=\frac{(1+\beta)\left(\frac{1}{1+2\beta}\right)^{\frac{1}{2\beta}}}{\beta+(1+\beta)^{-1-\frac{1}{\beta}}}>91.4\%$
of the maximum revenue. The ratio approaches 1 as $\beta\to 0$ (i.e., $F$ is sufficiently concave) or $\beta\to\infty$ (i.e., $F$ is sufficiently convex).
3.2 Welfare Implications
In this subsection, I study the welfare consequences of the number of positional good levels and exclusion in the revenue-maximizing mechanism. In particular, do full separation and exclusion also benefit consumers?
The consumer surplus (or agent welfare) is given by
$W=\int_{\theta_{0}}^{\bar{\theta}}U(\theta)\mathop{}\!\mathrm{d}F(\theta)=\int_{\theta_{0}}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)(s(\theta)+v^{\prime}(\theta))+U(\theta_{0})\mathop{}\!\mathrm{d}F(\theta)$ (11)
When the seller sets the price to maximize revenue, $U(\theta_{0})=0$ is pinned down by (IR). The following proposition provides comparative statics regarding the welfare consequences of (i) the number of positional good levels the seller offers and (ii) consumer participation.
Say $F$ satisfies IFR (DFR) if the failure rate $\frac{f(\theta)}{1-F(\theta)}$ is increasing (decreasing).
Proposition 4.
Assume $F$ satisfies IFR. Then,
  1. (i)
    Fixing exclusion $\theta_{0}$, 202020 The exclusion level is fixed when the price of the lowest-level good is fixed. In particular, when there is a free regular service, there is no exclusion (see Appendix A.2). the consumer surplus is decreasing in the number of positional good levels the seller offers.
  2. (ii)
    Increasing participation through a lower price increases the consumer surplus.
Assume $F$ satisfies DFR and Myerson’s regularity. Then, fixing exclusion $\theta_{0}$, the consumer surplus is increasing in the number of positional good levels the seller offers.
This proposition highlights the tension between the seller’s revenue and the consumer surplus: fixing consumer participation, offering more levels of positional goods increases the seller’s revenue but decreases the consumer surplus under IFR. Therefore, under IFR and when there is no exclusion, the regulator can increase consumer surplus by limiting the number of positional good levels (e.g., priority levels) the seller can offer.
Moreover, under IFR, expanding coverage by lowering the price increases the consumer surplus. Expanding coverage benefits the additional participants but also reduces the status of existing participants by pooling them with the new participants. For example, when the price of the regular service decreases, customers of the regular service will need to be pooled with more customers. Proposition 4 shows that under IFR, the marginal gain of additional participants outweighs the inframarginal loss of the existing participants.
A caveat is that if exclusion is not held fixed, it becomes ambiguous whether consumers suffer from more levels of positional goods offered by the seller. This is because when the seller offers more levels of positional goods, consumer participation also increases, which in turn increases the consumer surplus under IFR.
It is also worth noting that under DFR, if the distribution still satisfies Myerson’s regularity, the seller’s revenue and the consumer surplus are aligned—both are higher when the seller offers more levels of positional goods.
Consumer Surplus Maximization.
Relatedly, I study the optimal mechanism that maximizes consumer surplus. After rewriting the consumer surplus as
$W=\int_{\theta_{0}}^{\bar{\theta}}U(\theta)\mathop{}\!\mathrm{d}F(\theta)=\int_{\theta_{0}}^{\bar{\theta}}(\theta s(\theta)-p(\theta)+v(\theta))\mathop{}\!\mathrm{d}F(\theta),$ (12)
it is straightforward to show that when negative transfers are allowed and subject to budget balance ($\operatorname{\mathbf{E}}[p(\theta)]=0$), the consumer surplus is maximized by full separation without exclusion. Intuitively, this is because assortative matching is efficient, and transfers can redistribute efficiency gains among consumers.212121 See also Gershkov and Schweinzer (2010). I thank Alex Gershkov for pointing this out.
Observation 1.
When transfers are subject to budget balance, the consumer surplus-maximizing mechanism is $s(\theta)=F(\theta)$ and $p(\theta)=\int_{0}^{\theta}x\mathop{}\!\mathrm{d}F(x)-\operatorname{\mathbf{E}}[(1-F(\theta))\theta]$.
Nonnegative Prices.
Negative prices are implausible in some applications. In particular, in the applications to organizational hierarchies and education, the price (i.e., effort) must be nonnegative. Because the nonnegative price constraint also pins down $U(\theta_{0})$ through (IR), the characterization of the consumer surplus-maximizing mechanism under the constraint is similar to Proposition 4.
Proposition 5.
Under the nonnegative price constraint, the consumer surplus is maximized by
  1. (i)
    total pooling without exclusion (i.e., $s(\theta)=1/2$) and $p(\theta)=0$ if $F$ satisfies IFR;
  2. (ii)
    full separation without exclusion (i.e., $s(\theta)=F(\theta)$) and $p(\theta)=\int_{0}^{\theta}x\mathop{}\!\mathrm{d}F(x)$ if $F$ satisfies DFR.
Remark 10.
Under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$, the first (second) part still holds if $\gamma\in[1/2,1]$ ($\gamma\in[0,1/2]$). In particular, full separation is always optimal if $\gamma=0$ (Immorlica et al., 2015), while total pooling is always optimal if $\gamma=1$ (Frank, 1985b).
Remark 11.
If the lowest type is changed to $\underline{\theta}>0$, then some pooling at the bottom can increase consumer surplus even under DFR, as pooling at the bottom increases $U(\underline{\theta})=\underline{\theta}s(\underline{\theta})+v(\underline{\theta})>0$ by raising the lowest status $s(\underline{\theta})$.
The result shows that under the nonnegative price constraint, exclusion is detrimental to the consumer under either IFR or DFR. Under IFR, because it is optimal to pool participants, exclusion can be beneficial as it increases the status of the participants who would otherwise be pooled with the excluded consumers. The result shows that the marginal loss of the excluded consumers exceeds the inframarginal gain of participants, so exclusion is detrimental to the consumer. On the other hand, under DFR, exclusion has no benefits to participants because the optimal mechanism fully separate participants, so no exclusion is optimal.
The result implies that under IFR, the consumer-optimal regulation of positional goods without using subsidies is to restrict the seller to offer one level of positional good and serve all consumers (e.g., by providing a free regular boarding service).
The results also have several implications in organizational hierarchies and education, where the effort (price) has to be positive. On the one hand, if the type distribution satisfies the IFR property (i.e, has a thin tail), which is satisfied by many common distributions, completely eliminating status competitions (e.g., educational disarmament) is optimal for agent welfare. In other words, agents pay nothing and are pooled at the same status. In the education application, this implies complete educational disarmament—for example, random assignment in allocating slots at sought-after schools. On the other hand, if there are a few very high types (i.e., a heavy tail), such as the Pareto distribution, intense competition or meritocracy is optimal even from the agent’s perspective.
3.3 Social Planner
Assume that the social planner maximizes a weighted sum of the seller’s revenue and consumer surplus, so social welfare is $S=\int_{0}^{\bar{\theta}}\lambda p(\theta)+U(\theta)\mathop{}\!\mathrm{d}F(\theta)$, where $\lambda\geq 0$. In the monopolist provision problem where $p$ is transferred to the seller, $\lambda$ captures the weight of the seller’s revenue relative to the consumer surplus. When the price is allowed to be negative, I assume $\lambda\geq 1$, because otherwise the social planner would set $p(\theta)=-\infty$ to transfer all money from the seller to consumers.
In the education or the organization application, where $p$ is the agent’s nonnegative effort (so utilities are nontransferable), $\lambda$ captures the extent to which effort is socially productive. Agent effort is completely unproductive when $\lambda=0$ and productive when $\lambda=1$. When $\lambda>1$, the social benefit of effort is higher than the private benefit.
Denote $J_{\lambda}(\theta)=\lambda\theta-(\lambda-1)\frac{1-F(\theta)}{f(\theta)}$. By previous results, social welfare is
$S=\int_{\theta_{0}}^{\bar{\theta}}J_{\lambda}(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{0}}^{\bar{\theta}}\left(\lambda v(\theta)-(\lambda v^{\prime}(\theta)-1)\frac{1-F(\theta)}{f(\theta)}-(\lambda-1)U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta).$ (13)
Using the same argument as before, we have the following result.
Proposition 6.
The social welfare-maximizing mechanism fully separates the participants and potentially exclude some types
  1. (i)
    without the nonnegative price constraint and $\lambda\geq 1$, or
  2. (ii)
    under the nonnegative price constraint and if $J_{\lambda}(\theta)$ is increasing.
In general, the social welfare-maximizing mechanism depends crucially on the welfare weight $\lambda\geq 0$ on the seller relative to consumers, or the extent to which effort is productive relative to signaling in the application to education.
Corollary 6.1.
If $\lambda=1$, the social welfare-maximizing mechanism induces full separation without exclusion.
Intuitively, when $\lambda=1$, transfers are welfare-neutral. Hence, a finer partition strictly increases social welfare because it increases efficiency by matching the higher types (who value status more) with higher status.
4 Extensions
4.1 Intrinsic Quality
In this subsection, I consider the case where the positional good has an intrinsic quality ${q}$, produced at a cost $c({q})$, which determined its intrinsic value $v({q},\theta)=\theta q$ (à la Mussa and Rosen (1978)). Assume $c({q})$ is strictly increasing and strictly convex. Denote by $Q(\theta)=\operatorname{\mathbf{E}}_{\omega}[\chi(\theta,\omega)]$ the interim allocation. Thus, the buyer’s utility is $U(\theta)=\theta(s(\theta)+Q(\theta))-p(\theta)$.
Lemma 3.
A direct mechanism $\{p(\theta),s(\theta),Q(\theta)\}$ is incentive-compatible if and only if
  • $U(\theta)=U(0)+\int_{0}^{\theta}\left(s(x)+Q(x)\right)\mathop{}\!\mathrm{d}x$ for all $\theta\in[0,\bar{\theta}]$, and
  • $s(\theta)+Q(\theta)$ is increasing.
Using the same arguments as before, the revenue maximization problem is
$\max_{s\in\operatorname{\text{MPS}}(F),Q}\int_{0}^{\bar{\theta}}[J(\theta)(s(\theta)+Q(\theta))-c(Q(\theta))]\mathop{}\!\mathrm{d}F(\theta)$ (14)
subject to $s(\theta)+Q(\theta)$ is increasing. After showing that this monotonicity condition is equivalent to $s(\theta)$ and $Q(\theta)$ being increasing, the following proposition follows from the same arguments as Proposition 2.
Proposition 7.
Assume $J(\theta)$ is increasing. The revenue-maximizing mechanism excludes types below $\theta^{*}_{0}=J^{-1}(0)$ and offers $\chi^{*}(\theta)=c^{\prime-1}(J(\theta))$ to $\theta\geq\theta^{*}_{0}$.
The proposition implies that the optimal mechanism in revenue maximization is robust to the introduction of intrinsic quality. It also has implications for selling positional goods with intrinsic values. By attaching the status value, the seller can extract $\int_{\theta^{*}_{0}}^{\bar{\theta}}J(\theta)s^{*}(\theta)\mathop{}\!\mathrm{d}F(\theta)>0$ more revenue. For the uniform distribution on $[0,1]$ and quadratic cost $c(q)=q^{2}/2$, the revenue with status value is \$0.29 compared to \$0.08 with pure intrinsic value. This explains the amount of money luxury companies spent on marketing (to create status value) instead of research and development (to reduce the cost).
4.2 Deliberately Lowering the Status
In some applications, the principal can deliberately lower the status. In the priority service example (Example 2.1), the seller can prolong the waiting time $t=1-s$ by pausing service or procrastinating. Thus, the seller can create negative status $s<0$ or excessive waiting time $t>1$. 222222 For example, in the U.S., OPT processing time is excessively long for the regular service but takes only 30 days for the premium service.
In these cases, the range of status allocation $s$ becomes $[-M,1]$, where $M$ is a sufficiently large bound on the negative status (or the maximum excessive waiting time). Incentive compatibility still requires $s$ to be increasing, and feasibility requires $s\leq\hat{s}$ for some $\hat{s}\in\operatorname{\text{MPS}}_{w}(F)$. By lowering the status, the seller can offer negative status $s<0$ to low types and reduces their payoffs. Therefore, negative status (or excessive waiting time) serves as a substitute for exclusion.
Can the seller benefit from lowering the status or prolonging the waiting time? Intuitively, instead of excluding low-type agents, the seller can serve them with excessive waiting time, thereby extracting some revenue from them while preventing higher types from mimicking them. As shown in the following proposition, it is optimal to assign $s(\theta)=-v^{\prime}(\theta)$ to buyers with negative virtual valuation (i.e., $J(\theta)<0$). At the same time, under IFR, this increases the number of consumers who get zero payoffs (who are essentially excluded), so it also decreases the consumer surplus.
Proposition 8.
Assume $J(\theta)$ is increasing and that lowering the status is feasible. Then, the revenue-maximizing mechanism fully separates $\theta\geq J^{-1}(0)$ and assigns $s^{*}(\theta)=-v^{\prime}(\theta)$ to $\theta<J^{-1}(0)$.
Under IFR, allowing the seller to lower the status increases the seller’s revenue and decreases the consumer surplus. 232323 It strictly increases the seller’s revenue if $v$ is nonlinear (i.e., strictly concave or $v(0)>0$) and strictly decreases the consumer surplus if $v$ is nonlinear and $f(0)<\infty$.
Remark 12.
The allocation $s^{*}(\theta)=-v^{\prime}(\theta)$ is increasing because $v^{\prime\prime}(\theta)\leq 0$ and feasible because it can be induced by randomization.
Example 4.1 (Excessive Waiting Time).
Consider the payoff $u(p,s,\theta)=\theta s-p+v(\theta)$ ($=\tilde{v}(\theta)-\theta t-p$), where $t=1-s$ is the waiting time. Assume $\theta\sim\text{Unif }[0,1]$ and that $v(\theta)=v_{0}+\alpha\theta$, where $v_{0},\alpha>0$. Then, the revenue-maximizing mechanism is
$t^{*}(\theta)=\begin{cases}1-\theta,&\text{if }\theta\geq 1/2\\ 1+\alpha,&\text{if }\theta<1/2.\end{cases}$
Excessive waiting time $1+\alpha$ can be implemented by pausing service for $\alpha+1/4$ units of time before serving all types below $1/2$ in random order (i.e., pooling them into the same priority level).
Figure 3: Excessive Waiting Time
The insight that negative status serves as a substitute for exclusion also applies to the case where exclusion is impossible as the lowest-level good or service is free (see Appendix A). In this case, negative status or excessive waiting time diminishes the value of the free service and effectively plays the role of exclusion, thereby increasing the seller’s revenue. In particular, by offering a priority service and a free regular service, the seller extracts more revenue as she increases the waiting time for the free regular service.
4.3 Status Game As a Suffering
In this subsection, I consider the case where $v^{\prime}(\theta)\leq-1$, so the status game is a “suffering” in the sense that participants with higher types suffer more and receive lower payoffs. In the queuing example (Example 2.1), $v^{\prime}(\theta)\leq-1$ holds if higher types derive lower utility from the service after accounting for waiting costs—for example, if the service has the same value $v_{0}$ to everyone: $u(p,s,\theta)=v_{0}-\theta t-p$, where $t=1-s$ is the waiting time (and thus $v(\theta)=v_{0}-\theta$). Because higher types face higher costs or disutility, agents have incentives to overreport their types to get lower waiting time.
Formally, because $U^{\prime}(\theta)=s(\theta)+v^{\prime}(\theta)\leq 0$ for all participants when $v^{\prime}(\theta)\leq-1$, higher types have lower payoffs, thereby creating incentives for overreporting, in contrast to underreporting in the main specification. Therefore, it is high types that are potentially excluded. Specifically, there exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates (i.e., $U(\theta)\geq 0$) if and only if $\theta\leq\theta_{0}$. The feasibility condition in terms of majorization remains to hold for all participants $\theta\in[0,\theta_{0}]$.
The following proposition shows the results for the case where $v^{\prime}(\theta)\leq-1$.
Proposition 9.
The following results hold if $v^{\prime}(\theta)\leq-1$:
  1. (i)
    Assume $L(\theta)=\theta+F(\theta)/f(\theta)$ is increasing. Then,
    1. (a)
      The revenue is increasing in the number of positional good levels.
    2. (b)
      The revenue-maximizing mechanism excludes $\theta>\theta_{0}^{*}$ and fully separates $\theta\leq\theta_{0}^{*}$, where the optimal cutoff is $\theta^{*}_{0}=\operatorname*{arg\,max}_{\theta_{0}}\int_{0}^{\theta_{0}}L(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})F(\theta_{0})$.
    3. (c)
      If $-v^{\prime}(\theta)/v(\theta)\leq f(\theta)/F(\theta)$ for all $\theta\in[0,\bar{\theta}]$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=\bar{\theta}$).
  2. (ii)
    The seller can obtain at least half the maximum revenue by selling one level of positional good.
  3. (iii)
    Under the nonnegative price constraint, consumer surplus is maximized by total pooling (full separation) without exclusion if the reverse failure rate $\frac{f(\theta)}{F(\theta)}$ is decreasing (increasing).
  4. (iv)
    Deliberately lowering the status cannot increase the seller’s revenue.
Parts (i)–(iii) are qualitatively similar to those in the main specification, except that the revenue-maximizing mechanism excludes high types instead of low types. For part (iv), when the seller assign negative status, she has to assign it to low types first because $s$ is increasing (by incentive compatibility). Hence, because $U^{\prime}(\theta)\leq 0$, negative status reduces the payoffs of all participants, and therefore cannot increase the seller’s revenue.
Countervailing incentives.
A more complicated case is when $v^{\prime}(\theta)\in(-1,0)$, so the status game is undesirable for some types and desirable for others. Because $U^{\prime}(\theta)=s(\theta)+v^{\prime}(\theta)$ where $v^{\prime}(\theta)\in(-1,0)$ and $s\colon[0,\bar{\theta}]\to[0,1]$ is increasing, $U(\theta)$ is U-shaped, and agents have countervailing incentives (see also Lewis and Sappington, 1989): overreporting types when they are low and underreporting types when they are high. Therefore, it is the intermediate types that are potentially excluded.
5 Conclusion
In this paper, I study the allocation of positional goods using a mechanism approach. Many goods are positional in the sense that allocating an item to one agent has externalities on others: moving up the position of one consumer would push down that of another, and the value of the good depreciates as more people have it. The model has dual implications in the monopolist provision of positional goods and the design of status categories within organizations. The results also shed light on the trade-offs involved in educational disarmament.
In terms of the monopolist problem, the seller can often extract more revenue by offering many levels of positional goods through their status values. At the extreme, an all-pay auction with reserve prices, where the more buyers pay, the higher positions they receive, can maximize the seller’s revenue. On the other hand, offering more levels of positional goods is often detrimental to consumers. Thus, it is usually beneficial for consumers to restrict the number of position levels the seller can offer, and the consumer-optimal regulation without using subsidies is to restrict the seller to only one levels of positional goods and serving everyone. Nevertheless, in cases where a few buyers have very high valuations (e.g., Pareto distribution), both the seller’s revenue and the consumer surplus are higher when more status levels are offered—their objectives are aligned.
In terms of the organizational design problem, the results suggest that the management practice of stack ranking and rat races in education maximizes agents’ effort, consistent with the common wisdom, if the ability distribution satisfies Myerson’s regularity. The exclusion of low types can increase revenue or effort. On the contrary, from the agent’s perspective, less intense status competition is beneficial if the ability distribution satisfies IFR (or has a thin tail). This explains the call for abolishing stack ranking and educational disarmament, such as lotteries for school admissions. Nevertheless, when a few agents have very high abilities (e.g., Pareto distribution), rat races can be better for agent welfare. Thus, the objectives to maximize agent effort and welfare can be aligned, which challenges the common wisdom that rat races are always detrimental to agents. In either case, exclusion always harms agent welfare, even though it can increase the status of participants. Moreover, when learning effort is productive, education arms races can increase social welfare.
It remains to be explored how competition among sellers affects the results. This can be interpreted as offering a different track of education that leads to different status levels (e.g., vocational versus academic education) or having another product that has a separate queue. In either case, consumers of one set of positional goods derive utility independently of those of the other, either because it is psychologically incomparable or physically separate. If the other seller offers the same menu as the original seller, an immediate result is that each seller will attract half of the buyers, and that competition can increase consumer surplus.242424 Di Stasio et al. (2016) find that education is less positional in countries with more developed vocational education systems (e.g., Germany). However, when sellers compete in mechanisms strategically, the problem remains an open question.
Appendix A Optimal Mechanisms without Exclusion
In this section, I assume exclusion is impossible and that the lowest-level positional good is free. For example, in Moldovanu, Sela, and Shi (2007), agents cannot be excluded and will at least get the lowest status in the organization. In the benchmark model of Gershkov and Winter (2023), buyers can use the regular (non-priority) service for free. Because exclusion is not allowed, the free low-level position, rather than opting out, becomes the relevant outside option. As noted in Remark 2, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(F)$. I will repeat the analysis in Section 3 and show that the results are qualitatively similar.
A.1 Revenue Maximization
The revenue maximization problem is given by
$\max_{s(\theta),p(\theta)}\int_{0}^{\bar{\theta}}p(\theta)\mathop{}\!\mathrm{d}F(\theta)$ (A.1)
subject to the following constraints on $\theta\in[0,\bar{\theta}]$
$U(\theta)-v(\theta)=\theta s(\theta)-p(\theta)\geq\theta\underline{s}$ (IR) (A.2)
$U(\theta)-v(\theta)=U(0)-v(0)+\int_{0}^{\theta}s(x)\mathop{}\!\mathrm{d}x$ (IC) (A.3)
$s(\theta)\mbox{ is increasing}$ (A.4)
$s\in\operatorname{\text{MPS}}(F)$ (MPS) (A.5)
where $\underline{s}=\min\{s(\theta)\}$ denotes the lowest status, which is endogenously determined by the status allocation. The (IR) constraint, which is equivalent to $\tilde{U}(\theta)\equiv U(\theta)-v(\theta)-\theta\underline{s}\geq 0$, can be reduced to $U(0)-v(0)\geq 0$ because $U^{\prime}(\theta)-v^{\prime}(\theta)-\underline{s}=s(\theta)-\underline{s}\geq 0$.
The revenue is given by
$R=\int_{0}^{\bar{\theta}}p(\theta)\mathop{}\!\mathrm{d}F(\theta)$ $=\int_{0}^{\bar{\theta}}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)s(\theta)\mathop{}\!\mathrm{d}F(\theta)-(U(0)-v(0)).$ (A.6)
By (IR), it is optimal to set $U(0)=v(0)$. Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. The revenue maximization problem is equivalent to
$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta).$ (A.7)
Theorem 4 (Fan-Lorentz) in KMS implies the following result.
Proposition A.1.
Offering more positional good levels (strictly) increases the seller’s revenue if and only if $J(\theta)$ is (strictly) increasing. Thus, full separation maximizes revenue if and only if $J(\theta)$ is increasing.
Remark A.1.
GW’s Proposition 8 provides a sufficient condition for the seller’s revenue to be strictly increasing in the number of priority classes: $F$ satisfies the IFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is decreasing).
Remark A.2.
Effort maximization in MSS’s model is the same as revenue maximization because they assume linear effort costs (and no exclusion). They provide a sufficient condition for full separation in Theorem 4: $F$ satisfies the IFR property.
Infinitely many positional good levels can be implemented by an all-pay auction, in which the more money a consumer pays, the higher status he receives.
Proposition A.2.
The revenue-maximizing mechanism always separates the highest types $\theta\in[\bar{\theta}-\varepsilon,\bar{\theta}]$ for some $\varepsilon>0$.
{proofsketch}
Because $J^{\prime}(\bar{\theta})=2>0$, $J(\theta)$ is increasing in a neighborhood of $\bar{\theta}$, so by Proposition 2 in KMS, types in this neighborhood are separated.
Remark A.3.
This “separation at the top” result is a continuous-type analog of MSS’s Theorem 3.
Proposition A.3.
The seller can obtain at least half the maximum revenue by offering one positional good level in addition to a free low-level position.
Proof.
Denote by $p$ the price of the high-level position. The cutoff type $\theta(p)$ indifferent between two levels is given by
$\theta(p)\frac{1+F(\theta(p))}{2}-p=\theta(p)\frac{F(\theta(p))}{2}\iff\theta(p)=2p.$
By offering one positional good level in addition to a free low-level position, the seller’s maximum revenue is
$R_{2}=\max_{p}p(1-F(2p))=\frac{1}{2}\max_{p}p(1-F(p)).$
Because $\operatorname{\text{MPS}}(F)\subseteq\mathcal{M}\equiv\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$, we have
$\max R=\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)<\max_{q\in\mathcal{M}}\int_{0}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)=\max_{p}p(1-F(p))=2R_{2}.$
The inequality is strict because an extreme point of $\mathcal{M}$ is not in $\operatorname{\text{MPS}}({F})$. ∎
This proposition extends Proposition 3 to the case where exclusion is impossible.
Graphical Illustration.
Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$. Figure 4 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$. Assume $v(\theta)=0$ for simplicity. In Figure 4, (i) the area under the revenue curve (blue area in the left panel) represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)$, (ii) the red area in the right panel represents the revenue from having two levels of positional goods—a higher level offered at price $F(\theta^{*}_{0})$ in addition to a lower level offered for free, and (iii) the entire box of either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem. It is straightforward that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
Figure 4: Revenue curve $R(\tau)$ for the uniform distribution
The approximation performs well for many common distributions, as shown in the following examples (with $v(\theta)=0$).
Example (Exponential distribution).
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, offering one positional good level (in addition to a free low level) can obtain $73.6\%$ of the maximum revenue.
Example (Uniform distribution).
If $F(\theta)=\theta$ on $[0,1]$, offering one positional good level (in addition to a free low level) can obtain $75\%$ of the maximum revenue.
Example (Power distribution).
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, offering one positional good level (in addition to a free low level) can obtain
$\frac{R_{2}}{\max R}=\begin{cases}\frac{(1+2\beta)(1+\beta)^{-1/\beta}}{2\beta}>72.1\%,&\text{ if }\beta\geq 1\\ \frac{\beta(1+2\beta)(1+\beta)^{-1/\beta}}{2\beta^{2}-(1+\beta)^{-\frac{\beta+1}{\beta(1-\beta)}}\left(2\beta^{2}+(1+\beta)^{-\frac{\beta}{1-\beta}}-(1+\beta)\right)}>\frac{3}{4},&\text{ if }0<\beta<1\end{cases}$
of the maximum revenue. The ratio approaches 1 as $\beta\to\infty$ (i.e., $F$ is sufficiently convex) or $\beta\to 0$ (i.e., $F$ is sufficiently concave).
A.2 Consumer Surplus
The consumer surplus is given by
$W=\int_{0}^{\bar{\theta}}U(\theta)\mathop{}\!\mathrm{d}F(\theta)=\int_{0}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\operatorname{\mathbf{E}}[v(\theta)]+U(0)-v(0).$ (A.8)
Proposition A.4.
For any increasing allocations $s,\hat{s}\colon[0,\bar{\theta}]\to[0,1]$ such that $\hat{s}\in\operatorname{\text{MPS}}(s)$,
  1. (i)
    $\hat{s}$ results in a higher consumer surplus if $F$ satisfies IFR;
  2. (ii)
    $\hat{s}$ results in a lower consumer surplus if and only if $F$ satisfies DFR.
Therefore, the consumer surplus decreases (increases) as the seller offers more levels of positional goods if $F$ satisfies IFR (DFR).
Corollary A.4.1.
The consumer surplus is the highest when the seller offers
  1. (i)
    one position level (i.e., $s(\theta)=1/2$) if $F$ satisfies IFR;
  2. (ii)
    full separation (i.e., $s(\theta)=F(\theta)$) if and only if $F$ satisfies DFR.
In the former case, the consumer surplus is $\operatorname{\mathbf{E}}[v(\theta)+\theta/2]$.
Remark A.4.
GW’s Proposition 7 shows the sufficiency of IFR property ($\frac{1-F(\theta)}{f(\theta)}$ is decreasing) for no priority service to be consumer surplus-maximizing.
Remark A.5.
If both $\theta-\frac{1-F(\theta)}{f(\theta)}$ and $\frac{1-F(\theta)}{f(\theta)}$ are increasing (e.g., exponential and Pareto distributions), a finer partition (in terms of majorization) can increase both the seller’s revenue and consumer surplus. Thus, full separation maximizes both the revenue and consumer surplus.
Proposition A.5.
The consumer surplus is the highest when the seller offers one position level if and only if
$\int_{0}^{\theta}\left(\frac{1-F(x)}{f(x)}-\operatorname{\mathbf{E}}[\theta]\right)\mathop{}\!\mathrm{d}F(x)\geq 0\mbox{ for all $\theta\in[0,\bar{\theta}]$.}$
Proof.
Denote $I(\theta)=\int_{0}^{\theta}\frac{1-F(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)$. Then, the condition in the proposition is equivalent to $I(\theta)\geq I(\bar{\theta})F(\theta)=\operatorname{\mathbf{E}}[\theta]F(\theta)$ (graphically, $I(\theta)$ lies above the line connecting $I(0)=0$ and $I(\bar{\theta})=\operatorname{\mathbf{E}}[\theta]$ in the quantile space). Therefore, $I(\bar{\theta})F(\theta)$, which corresponds to total pooling, is the convex hull of $I$ (i.e., the largest convex function that lies below $I$).252525 A piecewise affine function consisting of other points on $I(\theta)$, which corresponds to offering more positional good levels, lies above $I(\bar{\theta})F(\theta)$ and leads to lower consumer surplus. By Proposition 2 in KMS, this condition is necessary and sufficient for total pooling to be welfare-maximizing. ∎
Remark A.6.
The condition is equivalent to $\operatorname{\mathbf{E}}[\tilde{\theta}\mid\tilde{\theta}\geq\theta]-\theta\leq\operatorname{\mathbf{E}}[\theta]$ for all $\theta\in[0,\bar{\theta}]$.
Remark A.7.
The condition is necessary and sufficient for customers’ welfare to be higher when one level of positional goods is offered than when any $k>1$ levels are offered. GW’s Proposition 1 provides a stronger sufficient condition for customers’ welfare to be higher when one level of positional good is offered than when two levels are offered. An even stronger sufficient condition is the IFR property, according to Proposition A.4.
Moreover, fix a type $\theta\in[0,\theta]$, maximizing his utility $U(\theta)$ subject to constraints (A.2)–(A.5) is equivalent to
$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)+U(0)-v(0).$ (A.9)
Because $p(0)\geq 0$, it is optimal to set $p(0)=0$ and $U(0)=v(0)$. The following proposition provides a necessary and sufficient condition for total pooling to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
Proposition A.6.
Every consumer’s utility is the highest when the seller offers one position level if and only if $F(\theta)\leq\theta/\bar{\theta}$ for all $\theta\in[0,\bar{\theta}]$ (i.e., $F$ first-order stochastic dominates the uniform distribution). A sufficient condition is that $f(\theta)$ is increasing.
{proofsketch}
$U(\theta)=\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)$. If $f$ is increasing, then $1/f$ is decreasing, so $s=1/2$ maximizes $U(\theta)$.
Note that $\int_{0}^{\theta}1/f(x)\mathop{}\!\mathrm{d}F(x)=\theta$, and the condition $F(\theta)\leq\theta/\bar{\theta}$ is equivalent to $\int_{0}^{\theta}(1/f(x)-\bar{\theta})\mathop{}\!\mathrm{d}F(x)\geq 0$ for all $\theta\in[0,\bar{\theta}]$. Similar to the proof of Proposition A.5, one can show that this condition is necessary and sufficient for $s(\theta)=1/2$ to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
Remark A.8.
Proposition A.6 provides the necessary and sufficient condition for all consumers to be worse off after any $k>1$ levels of positional goods are offered than if one level is offered (for free). GW’s Proposition 2 shows that if $F(c)\leq c/\bar{c}$, all consumers are worse off after the introduction of one additional level of priority service.
Appendix B Proofs
B.1 Proof of Section 3
Proof of Proposition 2.
The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to seller’s revenue given $\theta_{0}$, that is,
$R(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)+v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)\right)\mathop{}\!\mathrm{d}F(\theta).$ (B.1)
Lemma B.1.
Assume $J(\theta)$ is increasing. Fix a cutoff type $\theta_{0}$. For any two increasing allocations $s,\hat{s}\colon[\theta_{0},\bar{\theta}]\to[0,1]$ such that $\hat{s}\in\operatorname{\text{MPS}}(s)$, $s$ results in a higher revenue than $\hat{s}$.
Then, part (i) follows from Lemma B.1 because the allocation can be finer (and thus larger in the majorization order) as the number of positional good levels increases. Thus, the revenue-maximizing mechanism fully separates the participants (i.e., $s(\theta)=F(\theta)$) and the seller’s revenue is maximized by the cutoff
$\theta_{0}^{*}$ $=\operatorname*{arg\,max}_{\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)F(\theta)+v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)$
$=\operatorname*{arg\,max}_{\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$
Note that under the specification $\tilde{S}({q},G_{\chi})=\phi(S({q},G_{\chi}))$, where $\phi$ is an increasing and convex function, $J(\theta)s(\theta)$ in equation (B.1) is replaced by $J(\theta)\phi(s(\theta))$. Lemma B.1 still follows from Fan-Lorentz Theorem because $J(\theta)\phi(s)$ is convex in $s$ and supermodular in $(s,\theta)$ when $\phi$ is increasing and convex.
On the other hand, if $J(\theta)$ is decreasing on $\Theta_{p}\subseteq[0,\bar{\theta}]$, Fan-Lorentz Theorem implies that the revenue-maximizing allocation pools all types $\theta\in\Theta_{p}$ at the same level. ∎
Proof of Corollary 2.1.
Denote $J_{v}(\theta)=v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)$. Under IFR, $J_{v}(\theta)$ is increasing because $v(\theta),v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$. Denote
$\psi(\theta)\equiv-R^{\prime}(\theta)/f(\theta)=J(\theta)F(\theta)+J_{v}(\theta).$ (B.2)
Denote $\theta_{1}=J^{-1}(0)$. We have
$\psi(\theta_{1})=J(\theta_{1})F(\theta_{1})+J_{v}(\theta_{1})=v(\theta_{1})-\theta_{1}v^{\prime}(\theta)\geq 0$ (B.3)
by the concavity of $v(\theta)$ and $v(0)=0$. Moreover,
$\psi^{\prime}(\theta)=J^{\prime}(\theta)F(\theta)+J(\theta)f(\theta)+J_{v}^{\prime}(\theta)>0\mbox{ for all $\theta>\theta_{1}$.}$ (B.4)
Thus, for all $\theta>\theta_{1}$, we have $\psi(\theta)>0$ and thus $R^{\prime}(\theta)<0$. Hence, we must have $\theta^{*}_{0}\leq\theta_{1}$.
Moreover, because $J(\theta)$ is increasing, $J(\theta)F(\theta)>J(0)F(1)=-1/f(0)$. Thus, if $J_{v}(0)-1/f(0)=v(0)-(v^{\prime}(0)+1)/f(0)\geq 0$, the integrand is always positive, so no exclusion is optimal (i.e., $\theta_{0}^{*}=0$). ∎
Proof of Proposition 3.
Denoted by $p$ the price of the positional good. The cutoff type $\theta(p)$ indifferent between buying and not buying is given by
$\theta(p)\frac{1+F(\theta(p))}{2}-p+v(\theta(p))=0\implies p=\theta(p)\frac{1+F(\theta(p))}{2}+v(\theta(p)).$
Because $v^{\prime}(\theta)\geq 0$, all types above $\theta(p)$ will buy the good. Thus, by selling one level of positional good, the seller’s maximum revenue is
$R_{1}=\max_{\theta}\left(\theta\frac{1+F(\theta)}{2}+v(\theta)\right)(1-F(\theta))\geq\frac{1}{2}\max_{\theta}(\theta+v(\theta))(1-F(\theta))$
because $\frac{1+F(\theta)}{2}\theta\geq\frac{1}{2}\theta$. Assume first for simplicity $v(\theta)=0$.
Consider the auxiliary screening problem of selling an indivisible item of value to one buyer, in which a standard result implies a posted-price mechanism is optimal (see Börgers, 2015, Proposition 2.5). Formally, denote by $q$ the probability the item is allocated to $\theta$, and let =$\mathcal{M}=\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$. The agent’s payoff is $\theta q-p$. Then, the standard result implies that $\max_{q\in\mathcal{M}}\int_{0}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)=\max_{p}p(1-F(p))$, and any maximizer $q^{*}$ must be an extreme point of $\mathcal{M}$ (i.e., a posted-price mechanism).
Because $\operatorname{\text{MPS}}_{w}(F)\subseteq{\mathcal{M}}$, we have
$\max R=\max_{s\in\operatorname{\text{MPS}}_{w}(F)}\int_{0}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)<\max_{q\in{\mathcal{M}}}\int_{0}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)=\max_{p}p(1-F(p))\leq 2R_{1}.$
The inequality is strict because an extreme point of $\mathcal{M}$ is not in $\operatorname{\text{MPS}}_{w}({F})$ (except when everyone is excluded, which cannot be optimal).
Now consider $v(\theta)\geq 0$. In the auxiliary problem, the agent’s payoff becomes $(\theta+v(\theta))q-p$. We still have
$\max R$ $=\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$
$<\max_{q\in{\mathcal{M}},\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$
$=\max_{\theta}(\theta+v(\theta))(1-F(\theta))\leq 2R_{1}.$
In other words, an intrinsic value $v(\theta)\geq 0$ increases the maximum revenue from selling ordinary goods (which is smaller than $2R_{1}$) by at least the same amount as the maximum revenue from selling positional goods. ∎
Proof of Proposition 4.
(i) The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS.
Lemma B.2.
Fix a cutoff type $\theta_{0}$. For any two increasing allocations $s,\hat{s}\colon[\theta_{0},\bar{\theta}]\to[0,1]$ such that $\hat{s}\in\operatorname{\text{MPS}}(s)$, $\hat{s}$ results in a higher (lower) consumer surplus than $s$ if $F$ satisfies IFR (DFR).
Fix an exclusion level $\theta_{0}$. Denote by $s_{k}^{*}$ the revenue-maximizing status allocation chosen by the seller when she offers $k\geq 1$ levels of positional goods. Under Myerson’s regularity (which is implied by IFR), by Lemma B.1, $s_{k}^{*}\in\operatorname{\text{MPS}}(s_{k+1}^{*})$. Thus, by Lemma B.2, under IFR (DFR), the consumer surplus is decreasing (increasing) in $k$, the number of positional good levels the seller offers.
(ii) Fix any exclusion level $\theta_{0}$. If $s(\theta)$ is strictly increasing at $\theta_{0}$, then decreasing $\theta_{0}$ to $\theta_{0}-\varepsilon$ increases the consumer surplus of $\theta\in[\theta_{0}-\varepsilon,\theta_{0})$ and does not affect the consumer surplus of $\theta\in[\theta_{0},\bar{\theta}]$.
If $s(\theta)=s_{0}$ is constant on $[\theta_{0},\theta_{1}]$ for some $\theta_{1}\in(\theta_{0},\bar{\theta}]$ (and $s(\theta)>s_{0}$ for $\theta>\theta_{1}$), then the consumer surplus is
$W(\theta_{0})=\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+F(\theta_{1})/2+F(\theta_{0})/2\right)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{1}}^{\bar{\theta}}\frac{1-F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+s(\theta)\right)\mathop{}\!\mathrm{d}F(\theta),$
with derivative given by
[…]
Denote
$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta.$
Note that $z(\theta_{1})-F(\theta_{1})-v^{\prime}(\theta_{1})=0$, and under IFR,
$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}\leq\theta_{1}$. Hence, decreasing $\theta_{0}$ increases the consumer surplus $W(\theta_{0})$. ∎
Proof of Proposition 5.
(i) Under IFR, $\frac{1-F(\theta)}{f(\theta)}$ is decreasing, so the optimal $s^{*}(\theta)$ pools all types $\theta\in[\theta_{0},\bar{\theta}]$, i.e.,
$s^{*}(\theta)=(1+F(\theta_{0}))/2$ (B.5)
Now consumer surplus can be written as a function of the cutoff $\theta_{0}\in[0,\bar{\theta}]$:
$W(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+1/2+F(\theta_{0})/2\right)\mathop{}\!\mathrm{d}F(\theta)+v(0)\cdot\mathbf{1}[\theta_{0}=0]$ (B.6)
with derivatives on $\theta_{0}>0$ given by
[…]
It is straightforward that $W^{\prime}(\bar{\theta})=0$ and $W^{\prime\prime}(\bar{\theta})=(v^{\prime}(\theta_{0})+1)f(\bar{\theta})\geq 0$. Denote
$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+1/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta,$ (B.8)
Note that $z(\bar{\theta})-F(\bar{\theta})-v^{\prime}(\bar{\theta})=0$, and under IFR,
$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}>0$. Hence, no exclusion is optimal (i.e., $\theta^{*}_{0}=0$).
(ii) If $\frac{1-F(\theta)}{f(\theta)}$ is increasing, $s^{*}(\theta)=F(\theta)$ for all $\theta\geq\theta_{0}$, so
$W(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)(F(\theta)+v^{\prime}(\theta))\mathop{}\!\mathrm{d}F(\theta)+v(0)\cdot\mathbf{1}[\theta_{0}=0].$
Therefore, no exclusion is optimal because $W^{\prime}(\theta_{0})=-(1-F(\theta_{0}))(F(\theta_{0})+v^{\prime}(\theta))\leq 0$ for all $\theta_{0}>0$. ∎
B.2 Proof of Section 4
Proof of Proposition 8.
If $s(\theta)\geq-v^{\prime}(\theta)$, agents have incentives to underreport their types (because $U^{\prime}(\theta)>0$). Thus, it is optimal for the seller to assign the lowest status $s(\theta)=-v^{\prime}(\theta)$ for $\theta\leq J^{-1}(0)$. When $s(\theta)\leq-v^{\prime}(\theta)$, the negative status creates countervailing incentives for low-type buyers to overreport their types (because $U^{\prime}(\theta)<0$), and again it is optimal to assign $s(\theta)=-v^{\prime}(\theta)$ because $L(\theta)$ is always positive. Moreover, $p(\theta)=-\theta v^{\prime}(\theta)+v(\theta)\geq 0$ for all $\theta\leq J^{-1}(0)$ by the concavity of $v(\theta)$ and $v(0)\geq 0$. ∎
Proof of Proposition 9.
(i) After exchanging the order of integration, the revenue is
$R=\int_{0}^{\theta_{0}}\left(L(\theta)s(\theta)+v(\theta)+\frac{F(\theta)}{f(\theta)}v^{\prime}(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)$ (B.9)
(ii) The proof is similar to that of Proposition 3. The only difference is that because $v^{\prime}(\theta)\leq 0$, all types $\theta\leq\theta(p)$ buy the good. Thus, the auxiliary problem becomes a screening problem in procurement where the lower price get the good.
(iii) The proof is similar to that of Corollary 5. The only difference is that the consumer surplus is
$W=\int_{0}^{\theta_{0}}\left(-\frac{F(\theta)}{f(\theta)}(s(\theta)+v^{\prime}(\theta))\right)+U(\theta_{0})\mathop{}\!\mathrm{d}F(\theta).$ (B.10)
Thus, pooling (separating) all participants $\theta\in[\theta_{0},\bar{\theta}]$ is optimal if $\frac{F(\theta)}{f(\theta)}$ is increasing (decreasing).
In the case of pooling, $s(\theta_{0})=1-F(\theta_{0})/2$ for all $\theta\in[0,\theta_{0}]$.
$W(\theta_{0})=\int_{0}^{\theta_{0}}-\frac{F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+1-F(\theta_{0})/2\right)\mathop{}\!\mathrm{d}F(\theta)+v(\bar{\theta})\cdot\mathbf{1}[\theta_{0}=\bar{\theta}]$
Because $v^{\prime}(\theta_{0})\leq-1$, we have
$W^{\prime}(\theta_{0})=-{F(\theta_{0})}(v^{\prime}(\theta_{0})+1-F(\theta_{0})/2)+\frac{f(\theta_{0})}{2}\int_{0}^{\theta_{0}}F(\theta)\mathop{}\!\mathrm{d}\theta\geq 0$
for all $\theta_{0}<\bar{\theta}$. Hence, no exclusion is optimal (i.e., $\theta^{*}_{0}=\bar{\theta}$).
In the case of separation, we have $s(\theta)=F(\theta)$ for all $\theta\in[0,\theta_{0}]$. Thus, $\theta_{0}^{*}=\bar{\theta}$ because exclusion cannot benefit participants (and only harms the excluded buyers). Formally,
$W(\theta_{0})=\int_{0}^{\theta_{0}}-\frac{F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+F(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)+v(\bar{\theta})\cdot\mathbf{1}[\theta_{0}=\bar{\theta}]$
and $W^{\prime}(\theta_{0})=-{F(\theta)}\left(v^{\prime}(\theta)+F(\theta)\right)f(\theta)\geq 0$ for all $\theta_{0}<\bar{\theta}$. ∎
Proof of Proposition 7.
First, I show the monotonicity condition is equivalent to $s(\theta)$ or $Q(\theta)$ being increasing.
Lemma B.3.
$s(\theta)$ is (strictly) increasing if and only if $Q(\theta)$ is (strictly) increasing. Thus, $s(\theta)$ and $Q(\theta)$ are increasing if and only if $s(\theta)+Q(\theta)$ is increasing.
Proof.
Fix an allocation $\chi\colon\Theta\times\Omega\to\tilde{X}$. For any $\omega\in\Omega$ and $\theta_{1},\theta_{2}\in\Theta$, we have $S(\chi(\theta_{2},\omega),G_{\chi})>S(\chi(\theta_{1},\omega),G_{\chi})$ if and only if $\chi(\theta_{2},\omega)>\chi(\theta_{1},\omega)$ by the definition of $S$. Thus, $s(\theta)=\operatorname{\mathbf{E}}_{\omega}[S(\chi(\theta,\omega),G_{\chi})]$ is strictly increasing (constant) if and only if $Q(\theta)=\operatorname{\mathbf{E}}_{\omega}[\chi(\theta,\omega)]$ is strictly increasing (constant). ∎
Because the exclusion level $\theta^{*}_{0}=J^{-1}(0)$ in the standard case of pure intrinsic value coincides with that in the pure status value case in the previous section, the proposition follows directly from the lemma. ∎

### v3

@@
-Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments.
-The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”
+
 Abstract
-I study the optimal allocation of positional goods in the presence of externalities arising from consumers’ concerns for relative consumption.
-Using a mechanism design approach, I characterize the externalities by a feasibility condition.
-I show that the revenue-maximizing mechanism possibly excludes some buyers and fully separates the participants if their type distribution satisfies Myerson’s regularity.
-The seller can guarantee at least half the maximum revenue by offering a single level of positional goods.
-When there is no exclusion, as the seller offers more positional good levels, the consumer surplus decreases (increases) if the distribution has an increasing (decreasing) failure rate.
-Higher participation increases the consumer surplus under increasing failure rates.
+I study the optimal allocation of positional goods in the presence of externalities arising from consumers’ concerns about relative consumption.
 Applications include luxury goods, priority services, education, and organizational hierarchies.
-Keywords: Positional goods, mechanism design, externalities, feasibility.
+Using a mechanism design approach, I characterize the set of feasible allocations through a majorization condition.
+The revenue-maximizing mechanism possibly excludes some buyers and fully separates participants under Myerson’s regularity condition.
+The seller can guarantee at least half the maximum revenue by offering a single good.
+Without exclusion, offering more levels of goods decreases (increases) consumer surplus under increasing (decreasing) failure rates.
+Higher participation raises consumer surplus under increasing failure rates.
+Keywords: Positional goods, mechanism design, externalities.
 If everyone stands on tiptoe, no one sees better.
@@
 27).
-Because consumers care about their relative positions, allocating an item to one buyer has externalities on others.
+In essence, because consumers care about their relative positions, the allocation of positional goods creates externalities on others.
 Examples of positional goods abound.
-Consumer goods, especially luxury goods, are often positional because people derive happiness from comparisons (Frank, 1985a; Luttmer, 2005; Carlsson et al., 2007).
+Consumer goods are often positional because people derive happiness from comparisons (Frank, 1985a; Luttmer, 2005; Carlsson, Johansson-Stenman and Martinsson, 2007).
 Another example is priority services and front-running (Gershkov and Winter, 2023), such as priority boarding.
-Such services reduce consumers’ waiting time by moving them ahead of the crowd, but they become less effective as more consumers use them.
-Additionally, education is the classic example in Hirsch (1976): “The value to me of my education depends not only on how much I have but also on how much the man ahead of me in the job line has.”
-More broadly, status within an organization can also be viewed as a positional good, as employees invest efforts to move up in the hierarchy (Moldovanu, Sela, and Shi, 2007).
-As a revenue-maximizing seller, the organization designer can also induce effort through employees’ positional concerns and exclude low-ability employees.
-In these examples, the allocation of positional goods has externalities on consumers, as moving up the position of one consumer would push down that of another.
-At the same time, when more consumers buy the same good (e.g., priority boarding), its value depreciates due to more extensive use.
-Thus, unlike ordinary goods, it is impossible to allocate the good (e.g., the first place in line) to multiple consumers at face value: “What each of us can achieve, all cannot” (Hirsch, 1976).
-In general, positional goods induce a game of status, in which agents pay in money or effort to attain higher relative positions.
-111 Because of the dual interpretation of the model, I use “consumers” and “agents”, as well as “seller” and “principal” interchangeably.
-While the principal can exploit the status game to incentivize spending or effort, this often comes at the expense of agent welfare.
-222 For example, Asian governments have been orchestrating educational disarmament (The Economist, 2021).
+These services reduce consumers’ waiting time by moving them ahead of the crowd, but they become less effective as more consumers use them.
+Additionally, education is a classic example in Hirsch (1976): “The value to me of my education depends not only on how much I have but also on how much the man ahead of me in the job line has.” More broadly, status within an organization can be viewed as a positional good, as employees invest effort to move up in the hierarchy (Moldovanu, Sela and Shi, 2007).
+Similar to a revenue-maximizing seller, the organization designer induces effort through employees’ positional concerns and by excluding low-ability employees.
+In these examples, the allocation of positional goods has externalities on consumers.
+On the one hand, moving up the position of one consumer would push down that of another.
+On the other, when more consumers buy the same good, its value depreciates due to more extensive use.
+Therefore, unlike ordinary goods, it is impossible to allocate the same positional good (e.g., the first place in line) to multiple consumers without reducing its value: “What each of us can achieve, all cannot” (Hirsch, 1976).
+For this reason, positional goods remain scarce and competitive despite economic growth.
+In general, positional goods induce a game of status, in which agents pay money or exert effort to attain higher relative positions.
+[^1] Because the model has implications for both problems, I use “consumers” and “agents” interchangeably, as well as “seller” and “principal.” While the principal can exploit the status game to incentivize spending or effort, this often comes at the expense of agents’ welfare.
+[^2] For example, Asian governments have been orchestrating educational disarmament (The Economist, 2021).
 U.S.
@@
 In this paper, I study the optimal allocation of positional goods using a mechanism design approach.
-I assume buyers care about their relative positions, or status, which depends on the mass of consumers who buy goods at the same level and at lower levels.
+I assume buyers care about their relative positions, or status, which depends on the mass of consumers who buy goods at the same or lower levels.
 Moreover, they have private information about their types—i.e., valuations of the status (or marginal costs of effort).
 The seller can offer multiple levels of positional goods at different prices.
-The buyer can pay a price (or effort) for one of the goods to obtain the status value, or pay nothing and opt out.
-First, I characterize positional externalities by a feasibility condition: the interim status allocation is feasible (i.e., can be induced by an allocation of positional good) if and only if it is weakly majorized by the buyer’s type distribution in the quantile space.
+The buyer can pay a price (or exert effort) for one of the goods to attain the status, or pay nothing and opt out.
+First, I capture positional externalities using a feasibility condition: an interim status allocation is feasible (i.e., can be induced by an allocation of positional goods) if and only if it is weakly majorized by the buyers’ type distribution in the quantile space.
 Then, I focus on revenue maximization.
-Using results in Kleiner, Moldovanu, and Strack (2021, henceforth KMS), I show that offering more levels of positional goods increases the seller’s revenue if the type distribution is regular in the sense of Myerson (1981).
-Therefore, the revenue-maximizing mechanism excludes types with negative marginal revenue and fully separates participants, which can be implemented by an all-pay auction with a reserve price, where the buyer who pays more obtains the higher position.
-In contrast, if Myerson’s regularity condition fails for some types, the optimal mechanism pools them into the same level of positional goods.
-I also find that a simple posted-price mechanism that offers a single level of positional goods can guarantee at least half the maximal revenue.
-In the proof, I consider an auxiliary screening problem of selling a continuum of ordinary (i.e., nonpositional) goods to a continuum of buyers.
-The screening problem highlights the distinction between selling positional and ordinary goods: in the former, allocating the good to one buyer has externalities on others, so the seller cannot sell the highest status to multiple buyers at face value.
-Therefore, selling positional goods results in lower revenue compared to selling ordinary goods (Mussa and Rosen, 1978).
+Using results from Kleiner, Moldovanu and Strack (2021, henceforth KMS), I show that offering more levels of positional goods increases the seller’s revenue if the type distribution is regular in the sense of Myerson (1981).
+Therefore, the revenue-maximizing mechanism excludes types with negative marginal revenue and fully separates participants.
+This can be implemented by an all-pay auction with a reserve price, where the buyer who pays more obtains the higher position.
+In contrast, if Myerson’s regularity condition fails, the optimal mechanism will pool some types into the same level of positional goods.
+I also find that a posted-price mechanism that offers a single good at a fixed price can guarantee at least half the maximum revenue.
+For uniform and exponential distributions, the approximation ratio exceeds 90%.
+The approximation result is relevant when the levels of positional goods is subject to regulatory or practical constraints.
+It also highlights the role of positional externalities: In the standard problem of selling indivisible ordinary goods, where there are no positional externalities, a posted-price mechanism is revenue-maximizing.
+By contrast, when selling positional goods, the seller cannot offer the highest status without reducing its value to all buyers who pay the posted price.
+Consequently, a posted-price mechanism is no longer optimal, and the seller’s revenue is strictly lower.
 Next, I study how the number of positional good levels and exclusion affect consumer welfare.
@@
 On the other hand, when the type distribution has a decreasing failure rate (DFR), which implies a heavy tail, the consumer surplus is higher when the seller offers fewer positional good levels.
-Furthermore, under IFR, expanding consumer participation increases the consumer surplus, as the marginal gain of additional participants outweighs the inframarginal loss of the existing participants from reduced status due to pooling.
+Furthermore, under IFR, expanding consumer participation increases the consumer surplus, as the marginal gain of additional participants outweighs the inframarginal loss of the existing participants from the reduced status caused by pooling.
 I also study optimal mechanisms that maximize the consumer surplus.
 When negative transfers are allowed, full separation without exclusion maximizes the consumer surplus.
-Intuitively, this is because assortative matching, and possibly negative transfers can redistribute the efficiency gains among consumers (see also Gershkov and Schweinzer, 2010).
-By contrast, in applications where the price (effort) must be nonnegative, such as organizational hierarchies and education, I show that total pooling without exclusion maximizes the consumer surplus under IFR.
+Intuitively, this is because assortative matching is efficient, and possibly negative transfers can redistribute the efficiency gains among consumers and induce truth-telling (see also Gershkov and Schweinzer, 2010).
+By contrast, in applications where the price (i.e., effort) must be nonnegative, such as education and organizational hierarchies, I show that total pooling without exclusion maximizes the consumer surplus if the distribution satisfies IFR.
 Unlike selling ordinary goods, exclusion raises the status of participants who would be otherwise pooled with the excluded ones.
 Nevertheless, under IFR, the harm of exclusion to the excluded consumers exceeds its benefit to the participants.
-In contrast, if the distribution has a decreasing failure rate or a heavy tail (e.g., the Pareto distribution), the consumer surplus under the nonnegative price constraint is maximized by full separation without exclusion.
+In contrast, if the distribution has a heavy tail (e.g., the Pareto distribution), the consumer surplus under the nonnegative price constraint is maximized by full separation without exclusion.
+Intuitively, separation benefits high types but hurts low types, so it benefits agents on average when there are relatively more high types.
 For social welfare, if the social planner puts the same weight on the seller’s revenue and the consumer surplus, then full separation without exclusion is optimal because it leads to assortative matching, and transfer is welfare-neutral.
-As an extension, I consider the case where the intrinsic value of the positional good is determined by its intrinsic quality produced at a cost (à la Mussa and Rosen (1978)), and the results remain robust.
+As an extension, I show that my main results extend to the setting where the intrinsic value of the positional good is determined by its intrinsic quality produced at a cost (à la Mussa and Rosen (1978)).
 Moreover, I also allow the seller to deliberately lower the status, or equivalently, prolong the waiting time by pausing service or procrastinating.
 Thus, the seller can create negative status or excessive waiting time as a substitute for exclusion.
-I show that the revenue-maximizing mechanism fully separates those with positive virtual valuations and assigns negative status to those who with negative virtual valuations.
-Additionally, I consider cases where the status game is a “suffering” in that higher types suffer more from participation, and the results remain robust.
+I show that the revenue-maximizing mechanism fully separates those with positive virtual valuations and assigns a negative status to those with negative virtual valuations.
+Additionally, my results extend to the setting where the status game is a “suffering” in that higher types suffer more from participation.
 The model has dual interpretations that lead to two strands of applications.
 On the one hand, it is a monopolist problem of selling positional goods, such as priority services (Gershkov and Winter, 2023) and luxury goods.
-Alternatively, it can be interpreted as an organizational design problem where the principal motivates agent effort (e.g., Moldovanu, Sela, and Shi, 2007; Immorlica, Stoddard, and Syrgkanis, 2015) or maximizes agent welfare through status within an organization.
+Alternatively, it can be interpreted as an organizational design problem where the principal motivates agents’ effort (e.g., Moldovanu, Sela and Shi, 2007; Immorlica, Stoddard and Syrgkanis, 2015) or maximizes agents’ welfare through status within an organization.
 Under this interpretation, the optimal mechanism that maximizes effort possibly excludes some types and fully separates the participants if the ability distribution satisfies Myerson’s regularity.
 This can be seen in “rank-and-yank” organizations that rank every employee and terminate those who underperform.
-The implication also echoes the education example by Hirsch (1976) and sheds light on the consequences of education arms races.
-From the agent’s perspective, while the common wisdom calls for educational disarmament and less intense competition, I show that this benefits agents if their ability distribution has a thin tail.
-Under the IFR property, admission lotteries or random assignments, instead of meritocratic allocation, maximizes agent welfare.
-On the contrary, if the ability distribution has a heavy tail, meritocracy maximizes agents’ welfare (and their effort).
-Additionally, if education is sufficiently socially productive, educational arms races can be optimal for social welfare.
+The implications also echo the education example by Hirsch (1976) and shed light on the consequences of education arms races.
+From the agent’s perspective, while the common wisdom advocates educational disarmament and less intense competition, I show that this benefits agents if their ability distribution has a thin tail: Under the IFR property, admission lotteries or random assignments, instead of meritocratic allocation, maximizes agents’ welfare.
+On the contrary, if a few agents have very high abilities (i.e., a heavy tail), meritocracy maximizes agents’ welfare on average.
+Intuitively, low-ability agents benefit from soft competition, while high-ability agents prefer meritocracy, so on average, meritocracy raises (reduces) agents’ welfare when there are relatively more high-ability (low-ability) agents.
+Additionally, if education is sufficiently socially productive, meritocracy can be optimal for social welfare.
 In Appendix A, I consider the case where exclusion is impossible, which is true in many applications where the lowest status is free.
-For example, a society may not allow exclusion, so investing no effort guarantees the lowest social status (Moldovanu, Sela, and Shi, 2007).
+For example, a society may not allow exclusion, so investing no effort guarantees the lowest social status (Moldovanu, Sela and Shi, 2007).
 Moreover, priority services may be offered in addition to free regular services (Gershkov and Winter, 2023).
-In these cases, the lowest-level positional good may still have a positive value, as the lowest status cannot be offered to multiple agents at face value either.
-I provide the necessary and sufficient conditions for the revenue-maximizing mechanisms to be separating or pooling and for consumer welfare to be decreasing or increasing in the number of positional good levels, as well as some additional results, such as approximation by offering two levels.
+In these cases, the lowest-level positional good may still have a positive value, as the lowest status cannot be offered to multiple agents at its original value either.
+I provide necessary and sufficient conditions when the revenue-maximizing mechanism is separating or pooling, and when consumer welfare is decreasing or increasing in the number of positional good levels.
+I also show some additional results, such as the 2-approximation of the maximum revenue by selling a single good above the lowest level.
 The results are qualitatively similar to the benchmark model with exclusion.
@@
 The idea that the consumer’s utility depends on the comparison with others’ consumption dates back to Veblen (1899) (see also Duesenberry, 1949; Leibenstein, 1950; Bagwell and Bernheim, 1996).
-Since Hirsch (1976) introduced the concept of positional good, the status interdependence in preference has been formally modeled and studied by Frank (1985b, 2005, 2008), Robson (1992), and Hopkins and Kornienko (2004).333 Another strand of literature uses the Duesenberry formulation that stems from Duesenberry (1949) and Pollak (1976) (see the survey by Truyts (2010)), which assumes the consumer’s utility depends on the consumption of others in addition to their absolute consumption.
+Since Hirsch (1976) introduced the concept of positional good, the status interdependence in preference has been formally modeled and studied by Frank (1985b, 2005, 2008), Robson (1992), and Hopkins and Kornienko (2004).
+[^3] Another strand of literature uses the Duesenberry formulation that stems from Duesenberry (1949) and Pollak (1976) (see the survey by Truyts (2010)), which assumes the consumer’s utility depends on the consumption of others in addition to their absolute consumption.
 This literature focuses on consumer choice in the presence of positional externalities and the welfare effects of changes in the income distribution and government policies (e.g., luxury tax).
-Empirical evidence of positional goods and externalities include Frank (1985a, 2005), Luttmer (2005), Alpizar et al.
-(2005), Carlsson et al.
-(2007), Bursztyn et al.
+Empirical evidence of positional goods and externalities include Frank (1985a, 2005), Luttmer (2005), Alpizar, Carlsson and Johansson-Stenman (2005), Carlsson, Johansson-Stenman and Martinsson (2007), Bursztyn et al.
 (2018).
@@
 Durst (2021) finds empirical evidence that education is indeed positional in Germany.
-Moldovanu, Sela, and Shi (2007, henceforth MSS) study the optimal design of the number and size of status categories in an organization to maximize agents’ effort when they care about their relative position (i.e., status).
-They find that full separation is optimal if the ability distribution satisfies IFR, while coarse partition may be optimal depending on the distribution—for example, the optimal partition involves only two categories if it is sufficiently concave.
-Despite the similarities in our results, this paper is different in three ways.444 The definition of status in MSS, which equals the number of strictly inferior agents minus the number of strictly superior agents, is equivalent (up to an affine transformation) to my specification (see Footnote 10).
-First, I allow for exclusion of agents and stochastic allocations and study agent welfare.
-In reality, the exclusion of agents is common in organizations, and employee welfare is an important consideration.
-Second, in the case without exclusion, I obtain stronger results that are either necessary and sufficient conditions or require fewer distributional assumptions (see Appendix A).
-Third, I use a mechanism design approach with a feasibility characterization using majorization, while they use methods in optimal contest design (with finite agents).
+Moldovanu, Sela and Shi (2007, henceforth MSS) study the optimal design of the number and size of status categories in an organization to maximize agents’ effort when they care about their relative position (i.e., status).
+They find that full separation is optimal if the ability distribution satisfies IFR, while coarse partitions may be optimal for some distributions—for example, the optimal partition involves only two categories if the distribution is sufficiently concave.
+Because status within an organization is also a positional good, my framework generalizes theirs by allowing for exclusion of agents and stochastic allocations, while further analyzing agents’ welfare.
+[^4] The status definition in MSS—the number of strictly inferior agents minus the number of strictly superior agents—is equivalent (up to an affine transformation) to mine (see Footnote 10).
+In practice, organizations may exclude agents, and employee welfare is an important consideration.
+Methodologically, I use a mechanism design approach with a feasibility characterization using majorization, while they use methods in optimal contest design (with finite agents).
+[^5] In the case without exclusion, I obtain stronger results that are necessary and sufficient conditions or require fewer distributional assumptions (see Appendix A).
 Relatedly, Rayo (2013) studies a model of monopolist provision of conspicuous goods and characterizes the optimal allocation.
-His results are qualitatively similar to MSS’s, and he does not consider exclusion, consumer welfare maximization, or stochastic mechanisms either.
-Conspicuous goods in his definition also differ from positional goods (or status in MSS) in that consumers care about the average type in their assigned category instead of relative positions.
-Nevertheless, the feasibility condition can also be applied to conspicuous goods in his model (see Remark 4), which explains the similarities between his results and those of positional goods.555 Board (2009) assumes more general benefits from being in a category (i.e., group quality) and finds a contradictory result that the revenue-maximizing principal segregates agents too finely compared to the efficient level (which is full separation in my model).
-He also finds that revenue maximization leads to too much exclusion.
-Similar to MSS, Immorlica, Stoddard, and Syrgkanis (2015) study the optimal design of badges to incentivize users’ contribution (i.e., effort) to the online community through the status value.
-Unlike MSS, they consider exclusion and allow the status value to be concave or convex in the relative position.
-They find the “leaderboard with a cutoff” mechanism is optimal if the distribution satisfies Myerson’s regularity.
-As in MSS, they do not consider agent welfare or stochastic allocations, which is an important consideration in online communities.
-Most importantly, their status value function is strictly decreasing in the number of people weakly superior to them (without adjustment to those at the same level), which naturally precludes pooling from being optimal because it decreases the status of every agent pooled together.
-Consequently, (i) total pooling will never be optimal—even for agent welfare—in sharp contrast to my result, (ii) the optimality of full separation in effort maximization is less surprising, and (iii) exclusion is equivalent to pooling at the lowest status.
+His definition of conspicuous goods differs from positional goods in that consumers care about the average type in their assigned category instead of relative positions.
+Nevertheless, his results are qualitatively similar to MSS’s, and he also focuses on revenue maximization with deterministic mechanisms and without exclusion.
+In Section 4.4, I also show that the feasibility condition can also be applied to conspicuous goods in his model, which explains the similarities between our results for revenue maximization.
+[^6] Board (2009) assumes that the value of being in a group category takes a more general form and finds that the revenue-maximizing mechanism segregates agents too finely compared to the efficient level under a regularity condition.
+He also finds that revenue maximization excludes too many agents.
+Similar to MSS, Immorlica, Stoddard and Syrgkanis (2015) study the optimal design of badges to incentivize users’ contribution (i.e., effort) to the online community through the status value.
+Although they consider exclusion and allow the status value to be concave or convex in the relative position, they do not consider agents’ welfare or stochastic allocations either.
+Most importantly, their status value function is decreasing in the number of people weakly superior to them, without adjustment for those at the same level.
+This specification precludes pooling from being optimal because it reduces the status of all pooled agents.
+Consequently, total pooling is never optimal—even for agents’ welfare—in sharp contrast to my result, and exclusion is equivalent to pooling at the lowest status.
 In a different vein, Gershkov and Winter (2023, henceforth GW) study a model of priority service and analyze its implications for consumer welfare.
-Priority service (PS) is an example of positional goods that arise from technological constraints.
-Using a mechanism design approach, this paper generalizes their model by allowing for multiple (and stochastic) PS levels and exclusion.
-By contrast, they focus on two PS levels instead of the optimal mechanism.
-666 They extend their model to multiple PS levels but without exclusion in Section V.
+Because priority service (PS) is a positional good arising from technological constraints, my model extends theirs to multiple (and stochastic) PS levels while allowing for exclusion using a mechanism design approach.
+By contrast, they focus on the welfare implications of having two PS levels instead of the optimal mechanism.
+[^7] They extend their model to multiple PS levels, but without exclusion, in Section V.
 In the case without exclusion, I refine some of their results by providing necessary and sufficient conditions and show some additional results in Appendix A.
 Another closely related paper is by Loertscher and Muir (2022), who study revenue maximization when selling a fixed inventory of goods with heterogeneous qualities.
-Without the intrinsic value and under linear status, my model is equivalent to a continuous version of theirs where each quality has a unit stock.
+Because positional externalities can arise from a limited stock, under the linear status specification and without the intrinsic value, my model can be viewed as a continuous version of theirs where each quality has a unit stock.
 In particular, pooling multiple types into the same status corresponds to randomization (or conflation) in their model.
+Nevertheless, my modeling approach leads to various different applications in positional goods, such as education and organizational status.
+Moreover, because positional externalities may also arise from psychological reasons, I also consider nonlinear (concave or convex) status specifications and different magnitudes of concern about the number of consumers at the same level.
 More broadly, this paper is also related to mechanism design with allocative externalities (e.g.
-Jehiel, Moldovanu, and Stacchetti, 1996; Jehiel and Moldovanu, 2005; Dworczak, Reuter, Kominers, and Lee, 2024).
+Jehiel, Moldovanu and Stacchetti, 1996; Jehiel and Moldovanu, 2005; Csorba, 2008; Ostrizek and Sartori, 2023; Akbarpour, Dworczak and Kominers, 2024; Dworczak et al., 2024).
+In particular, this model focuses on a large population of agents, and the positional externalities arise from their relative consumption instead of types, group identities, or network effects.
 2 Model
@@
 A monopolist seller (she) sells positional goods $q\in X\subseteq\mathbb{R}_{++}$ to a continuum of buyers (he) with unit mass.
+A larger $q$ represents a higher-level good, such as a larger car or a higher boarding class.
 Denote by $q=0$ the outside option of not buying the good.
@@
 Buyers are heterogeneous in their valuations of status (i.e., types), denoted by $\theta$.
-Assume the buyer’s type $\theta$ has a distribution $F(\theta)$ with full support $\Theta=[0,\bar{\theta}]$ and density $f(\theta)$.777 The support can be extended to $\bar{\theta}=\infty$ and $\underline{\theta}>0$.
+Assume the buyer’s type $\theta$ has a distribution $F(\theta)$ with full support $\Theta=[0,\bar{\theta}]$ and density $f(\theta)$.
+[^8] The support can be extended to $\bar{\theta}=\infty$ and $\underline{\theta}>0$.
 The latter case leads to a complication in consumer surplus maximization under decreasing failure rates (see Remark 7).
-A buyer with type $\theta$ who pays $p$ for a status $s\in[0,1]$ from the good receives a payoff of $u(p,s,\theta)=\theta s-p+v(\theta)$, where $v(\theta)\geq 0$ is the intrinsic value of the good.
-888 I assume for now that the intrinsic value is independent of $q$ to abstract from quality differentiation (Mussa and Rosen, 1978).
-In Section 4.1, I consider an extension where the positional good also has an intrinsic quality equal to ${q}$ produced at a cost $c({q})$.
-Assume $v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$, so higher types have higher intrinsic values.
-999 In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$, so higher types have lower intrinsic values.
-If the buyer does not get the good, his status is defined as $s=0$, and he does not get the intrinsic value.
+A buyer with type $\theta$ who pays $p$ and acquires status $s\in[0,1]$ from the good receives a payoff of $u(p,s,\theta)=\theta s-p+v(\theta)$, where $v(\theta)\geq 0$ is the intrinsic value of the good.
+[^9] I assume for now that the intrinsic value is independent of $q$ to abstract from quality differentiation (Mussa and Rosen, 1978).
+In Section 4.1, I consider an extension where the positional good also has an intrinsic quality equal to q{q} produced at a cost c(q)c({q}).
+Assume $v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$, so that higher types have higher intrinsic values.
+If the buyer does not receive the good, his status is defined as $s=0$, and he does not obtain the intrinsic value.
 Definition of status.
 Denote by $G$ the distribution of buyers’ consumption of the good $q$ over $\tilde{X}\equiv X\cup\{0\}$.
-Then, $G({q})$ represents the mass of buyers who consume weakly lower-level goods than ${q}$ (e.g., smaller cars, lower boarding groups) or opt out ($q=0$).
+Then, G(q)G({q}) represents the mass of buyers who consume weakly lower-level goods than q{q} or opt out.
 Following Robson (1992) and Hopkins and Kornienko (2004), I define the status of a buyer who consume $q\in X$ as
 $S({q},G(\cdot))=\frac{G^{-}({q})+G({q})}{2}\in[0,1],$ (1)
-where $G^{-}({q})\equiv\lim_{{q}\to{q}^{-}}G({q})$ is the mass of buyers who consume strictly lower-level goods than ${q}$ (or opt out).
-In other words, the status equals the mass of buyers who consume strictly lower-level goods plus half the mass of buyers at the same level.101010 Hopkins and Kornienko (2004) assume a more general specification: $S(q,G(\cdot))=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1)$.
+where $G^{-}({q})\equiv\lim_{{q}\to{q}^{-}}G({q})$ is the mass of buyers who consume strictly lower-level goods than q{q} (or opt out).
+In other words, the status equals the mass of buyers who consume strictly lower-level goods plus half the mass of buyers at the same level.
+[^10] Hopkins and Kornienko (2004) assume a more general specification: $S(q,G(\cdot))=\gamma G({q})+(1-\gamma)G^{-}({q})$, where $\gamma\in[0,1)$ captures the concern about the number of consumers at the same level.
 If $\gamma>1/2$ ($\gamma<1/2$), pooling will decrease (increase) the total expected status.
@@
 Below is another example in queuing.
-Example 2.1 (Priority Services (Gershkov and Winter, 2023)).
-A unit mass of customers arrive at the same time.
-A seller can serve one customer at each instant and offers priority services to skip the line: consumers at higher priority levels are served first, and those of the same level are served in random order.
+Example 1 (Priority Services (Gershkov and Winter, 2023)).
+A unit mass of consumers arrive at the same time, and the seller can serve one at each instant.
+The seller provides priority services to skip the line: consumers at higher priority levels are served first, and those at the same level are served in random order.
 The consumer’s payoff is $u(p,t,\theta)=\tilde{v}(\theta)-\theta t-p$, where $t$ is the waiting time, $\theta$ is the value of time (or the cost of waiting), and $\tilde{v}(\theta)$ is the value of the service.
-This is equivalent to the model with $s=1-t$ and $v(\theta)=\tilde{v}(\theta)-\theta$.
-111111 The assumption $v^{\prime}(\theta)\geq 0$ requires $\tilde{v}^{\prime}(\theta)\geq 1$ here, that is, higher types derive higher utility from the service after accounting for waiting costs.
-In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$ (i.e., $\tilde{v}^{\prime}(\theta)\leq 0$) so that higher types derive higher utility from the service after accounting for waiting costs.
-The seller essentially sells the waiting time $t\in[0,1]$ subject to the feasibility constraint
+This is equivalent to the model with s=1ts=1-t and $v(\theta)=\tilde{v}(\theta)-\theta$.
+[^11] The assumption $v^{\prime}(\theta)\geq 0$ implies $\tilde{v}^{\prime}(\theta)\geq 1$ here, that is, higher types derive lower utility from the service after accounting for waiting costs.
+In Section 4.5, I consider an extension where $v^{\prime}(\theta)\leq-1$ (i.e., $\tilde{v}^{\prime}(\theta)\leq 0$) so that higher types derive higher utility from the service after accounting for the waiting cost.
+The priority services provider essentially sells the waiting time $t\in[0,1]$ subject to the feasibility constraint
 $t=\frac{\Phi^{-}({t})+\Phi({t})}{2}\in[0,1],$
 where $\Phi$ is the distribution of the waiting time $t$.
-Because the status category within organization is also a positional good, another example is the status contests in organizations, where the principal (seller) maximizes effort by designing status categories of agents (buyers).
-Example 2.2 (Status Contests (Moldovanu et al., 2007)).
-Consider a principal who maximizes agent effort by designing status categories.
-Agents care about their relative positions within the organization.
-Then, the positional good $q$ represents the status category, the price $p\geq 0$ represents the agent’s effort, and the type $\theta$ represents the agent’s ability (i.e., marginal cost of effort).
-The agent has a linear effort cost $p/\theta$ and a payoff $s-p/\theta$ if he has status $s\in[0,1]$ in the organization.
+Because the status category within organization is also a positional good, another example is the status contests in organizations, where the principal maximizes effort by designing status categories of agents.
+Example 2 (Status Contests (Moldovanu, Sela and Shi, 2007)).
+Consider a principal who maximizes agents’ effort by designing status categories when they care about their status within the organization.
+The positional good $q$ represents the status category, the price $p\geq 0$ represents the agent’s effort, and the type $\theta$ represents the agent’s ability (i.e., marginal cost of effort).
+The agent has a linear effort cost $p/\theta$ and receives a payoff $s-p/\theta$ if he attains status $s\in[0,1]$.
 When the agent is excluded or expelled, he receives zero payoff.
@@
 2.2 A Mechanism Design Approach
-Consider a direct mechanism $(p(\theta),\chi(\theta,\omega))$, consisting of a payment function $p\colon\Theta\to\mathbb{R}$ and a (potentially stochastic) positional good allocation $\chi\colon\Theta\times\Omega\to\tilde{X}=X\cup\{0\}$, where the random variable $\omega\in\Omega=[0,1]$ captures the randomness in the mechanism.121212 The allocation can be implemented by a pricing scheme as a function of the distribution of positional good levels.
+Consider a direct mechanism $(p(\theta),\chi(\theta,\omega))$, consisting of a payment function $p\colon\Theta\to\mathbb{R}$ and a (potentially stochastic) positional good allocation $\chi\colon\Theta\times\Omega\to\tilde{X}=X\cup\{0\}$, where the random variable $\omega\in\Omega=[0,1]$ captures the randomness in the mechanism.
+[^12] The allocation can be implemented by a pricing scheme as a function of the distribution of positional good levels.
 An allocation $\chi$, together with the type distribution $F$, induces a distribution of positional goods $q$, given by $G_{\chi}(q\mid\omega)=\int\mathbf{1}[\chi(\theta,\omega)\leq q]\mathop{}\!\mathrm{d}F(\theta)$.
@@
 In particular, under a deterministic allocation $\chi\colon\Theta\to\tilde{X}$, the interim status is $s(\theta)=S(\chi(\theta),G_{\chi})$.
-Denote by $U(\theta)=\theta s-p(\theta)+v(\theta)$ the buyer’s payoff if he participates and $U(\theta)=0$ otherwise.
-131313 Although stochastic mechanisms are allowed, it is never optimal to randomly exclude participants either for the seller or consumers.
+Denote by $U(\theta)=\theta s(\theta)-p(\theta)+v(\theta)$ the buyer’s payoff if he participates and $U(\theta)=0$ otherwise.
+[^13] Although stochastic mechanisms are allowed, it is never optimal to randomly exclude participants either for the seller or consumers.
 This can be formally shown by multiplying $U(\theta)$ (or $v(\theta)$) by an expected participation dummy $\delta(\theta)\in[0,1]$and showing that $\delta^{*}(\theta)=1$ for all $\theta\geq\theta_{0}$.
@@
 Lemma 1.
-There exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates (i.e., $U(\theta)\geq 0$) if and only if $\theta\geq\theta_{0}$.
+There exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates if and only if $\theta\geq\theta_{0}$.
 Lemma 2.
@@
 Formally, say an interim status $s(\theta)$ is feasible if there exists an allocation $\chi$ that induces it—i.e., $s(\theta)=\operatorname{\mathbf{E}}_{\omega}[S(\chi(\theta,\omega),G_{\chi}(\cdot\mid\omega))]$.
-To provide intuitions, consider a deterministic allocation without exclusion $\chi\colon\Theta\to{X}$.
+To provide an intuition, consider a deterministic allocation $\chi\colon\Theta\to\tilde{X}$.
 Incentive compatibility implies that $\chi(\theta)$ is increasing because $s(\theta)=S(\chi(\theta),G_{\chi})$ is increasing.
 When $\chi(\theta)$ is strictly increasing at $\theta$, then $s(\theta)=F(\theta)$.
-Otherwise, when $\chi(\theta)=q>0$ is constant on some interval $[\theta_{1},\theta_{2}]$ (and $\chi(\theta)\neq q$ outside this interval), then
+Otherwise, when $\chi(\theta)=q>0$ is constant on some interval $\theta_{1},\theta_{2}]$ (and $\chi(\theta)\neq q$ outside this interval), then
 $s(\theta)=\frac{G^{-}(q)+G(q)}{2}=\frac{F(\theta_{1})+F(\theta_{2})}{2}=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\tilde{\theta}\in[\theta_{1},\theta_{2}]]=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\chi(\tilde{\theta})=\chi(\theta)]$
 for all $\theta\in[\theta_{1},\theta_{2}]$. In either case, $s(\theta)=S(\chi(\theta),G_{\chi})=\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid\chi(\tilde{\theta})=\chi(\theta)]$.
-Now I characterize the necessary and sufficient condition for an increasing interim status $s(\theta)$ to be feasible in general.
+Now I provide the necessary and sufficient condition for an increasing interim status $s(\theta)$ to be feasible.
 Theorem 1 (Feasibility).
@@
 $\int_{x}^{\bar{\theta}}s(\theta)\mathop{}\!\mathrm{d}F(\theta)\leq\int_{x}^{\bar{\theta}}F(\theta)\mathop{}\!\mathrm{d}F(\theta)\text{ for all }x\in[0,\bar{\theta}].$ (3)
-The proof is à la Theorem 3 (Border’s Theorem) in KMS.
-Proof.
-($\implies$) follows from $\operatorname{\mathbf{E}}[s(\theta)\mid\theta\geq x]\leq\operatorname{\mathbf{E}}[F(\theta)\mid\theta\geq x]$ for all $x\in[0,\bar{\theta}]$ because switching from any $s(\theta)$ to assortative rule ($s(\theta)=F(\theta)$) takes the status of low type to high types.
-[…]
-$s(\theta)=\frac{1}{F(\bar{\theta}_{i})-F(\underline{\theta}_{i})}\int_{\underline{\theta}_{i}}^{\bar{\theta}_{i}}F(\theta)\mathop{}\!\mathrm{d}F(\theta)=\frac{F(\underline{\theta}_{i})+F(\bar{\theta}_{i})}{2};$
-and (iii) fully separating otherwise, where $s(\theta)=F(\theta)$.
-Hence, by Proposition 1 (Choquet’s Theorem) in KMS, measure $\lambda_{s}$ supported on the extreme points of $\operatorname{\text{MPS}}_{w}(F)$ such that $s=\operatorname{\mathbf{E}}[\tilde{s}\mid\tilde{s}\sim\lambda_{s}]$.
-Thus, $s$ can be induced by randomizing over allocations that each induces extreme points of $\operatorname{\text{MPS}}_{w}(F)$.
-∎
+The proof (in Appendix B) is à la Theorem 3 (Border’s Theorem) in KMS.
 Remark 1.
-For deterministic allocations, $s(\theta)$ must be an extreme point of $\operatorname{\text{MPS}}_{w}(F)$—i.e., either the majorization constraint or the monotonicity constraint binds on $[\theta_{0},\theta]$ and $s(\theta)=0$ for all $[0,\theta_{0}]$ (see Theorem 1 and Corollary 2 in KMS).
-For stochastic allocations, $s(\theta)$ can take other forms.
+For deterministic allocations $\chi\colon\Theta\to\tilde{X}$, $s(\theta)$ must be an extreme point of $\operatorname{\text{MPS}}_{w}(F)$—i.e., either the majorization constraint or the monotonicity constraint binds on $\theta_{0},\bar{\theta}]$ and $s(\theta)=0$ for all $0,\theta_{0}]$.
+For stochastic allocations $\chi$, $s(\theta)$ can take other forms.
 Remark 2.
 When exclusion is impossible, the equality must hold at $x=0$, that is, $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[F]=1/2$.
-In other words, $s(\theta)$ is a mean-preserving spread of $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}(F)$.141414 Without exclusion, one can also write $s(\theta)=\operatorname{\mathbf{E}}_{{q}\sim\chi(\theta,\Omega)}[\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid{q}]]$, where $\chi(\theta,\Omega)$ denotes the distribution of $q=\chi(\theta,\omega)$ given $\theta$ generated by the randomness $\omega\in\Omega$.
-Remark 3.
-Under MSS’s specification $S({q},G)=G^{-}({q})-(1-G({q}))$ without exclusion, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(2F-1)$ with $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[2F-1]=0$.
-Remark 4.
-In Rayo’s (2013) model of conspicuous goods, where the status value induced by $\chi$ is $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$, the feasibility condition is $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$.
-Intuitively, we have $s(\theta)=F(\theta)$ under full separation.
-In contrast, any (partial) pooling of multiple types into the same positional good makes the status $s(\theta)$ a mean-preserving spread of $F(\theta)$ in the quantile space because the value of the positional good becomes lower than its face value, or the value under full separation.
+Formally, $s(\theta)$ is a mean-preserving spread of $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}(F)$.
+[^14] Without exclusion, one can also write $s(\theta)=\operatorname{\mathbf{E}}_{{q}\sim\chi(\theta,\Omega)}[\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid{q}]]$, where $\chi(\theta,\Omega)$ denotes the distribution of $q=\chi(\theta,\omega)$ given $\theta$ generated by the randomness $\omega\in\Omega$.
+[^15] Under MSS’s specification $S({q},G)=G^{-}({q})-(1-G({q}))$ without exclusion, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(2F-1)$ with $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[2F-1]=0$.
+Intuitively, under full separation, the interim status equals the quantile (i.e., $s(\theta)=F(\theta)$).
+In contrast, (partially) pooling multiple types into the same positional good makes the status $s(\theta)$ a mean-preserving spread of $F(\theta)$ in the quantile space because the value of the positional good becomes lower than its face value—i.e., the value under full separation.
 Example (Two Status Levels).
@@
 $s(\theta)=\begin{cases}s_{L}=F(\theta^{*})/2,&\text{ if }\theta<\theta^{*},\\ s_{H}=(1+F(\theta^{*}))/2,&\text{ otherwise}.\end{cases}$
-It is straightforward to check that $\operatorname{\mathbf{E}}[s]=1/2$ and that $\int_{\theta}^{\bar{\theta}}s(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})\leq\int_{\theta}^{\bar{\theta}}F(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})$ for all $\theta\in[0,\bar{\theta}]$ (with equality at 00, $\theta^{*}$, and $\bar{\theta}$).
+It is straightforward to check that $\operatorname{\mathbf{E}}[s]=1/2$ and that $\int_{\theta}^{\bar{\theta}}s(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})\leq\int_{\theta}^{\bar{\theta}}F(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})$ for all $\theta\in[0,\bar{\theta}]$ (with equality at 0, $\theta^{*}$, and $\bar{\theta}$).
 Figure 1: Two Priority Levels
@@
 Two extreme mechanisms are particularly of interest: full separation (i.e., $s(\theta)=F(\theta)$) and total pooling among participants (i.e., $s(\theta)=(1+F(\theta_{0}))/2$).
-The former can be induced by offering a continuum of positional goods that lead to a proliferation of status (i.e., the finest partition), and the latter can be induced by offering only one level of positional good.
+The former can be induced by offering a continuum of positional goods that lead to a proliferation of status (i.e., the finest partition), and the latter can be induced by offering a single good.
 3.1 Revenue Maximization
-The revenue (or effort) maximization problem is given by
+The revenue maximization problem is given by
 $\max_{s(\theta),p(\theta),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}p(\theta)\mathop{}\!\mathrm{d}F(\theta)$ (4)
@@
 $s\in\operatorname{\text{MPS}}_{w}(F)$ (MPS) (8)
-Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$.
+Define $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$.
 Say $F$ satisfies the Myerson’s regularity if $J(\theta)$ is increasing.
 Say $F$ has an increasing failure rate (IFR) if $\frac{1-F(\theta)}{f(\theta)}$ is decreasing.
-The revenue is given by
+By standard arguments, expected revenue is given by
 $R=\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)-U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$ (9)
-[…]
-$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (10)
+By (IR), it is optimal to set $U(\theta_{0})=0$.
+[^16] The price is given by $p(\theta)=\theta_{0}s(\theta_{0})+v(\theta_{0})+\int_{\theta_{0}}^{\theta}x\mathop{}\!\mathrm{d}s(x)\geq 0$.
+Thus, the revenue maximization problem is equivalent to
+$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (10)
 The maximum is attained at the extreme point of $\operatorname{\text{MPS}}_{w}(F)$, which is an extreme point of $\operatorname{\text{MPS}}(F\cdot\mathbf{1}_{[\theta_{0},\bar{\theta}]})$.
-The following proposition follows from Fan-Lorentz Theorem (see, e.g., Theorem 4 in KMS).
+The following proposition follows from the Fan-Lorentz Theorem (see, e.g., Theorem 4 in KMS).
 Proposition 2.
@@
     The revenue-maximizing mechanism excludes $\theta<\theta_{0}^{*}$ and fully separates $\theta\geq\theta_{0}^{*}$, where the optimal cutoff is $\theta^{*}_{0}=\operatorname*{arg\,max}_{\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0}))$.
-If $J(\theta)$ is decreasing on some $\Theta_{p}\subseteq[0,\bar{\theta}]$, then the revenue-maximizing mechanism pools every $\theta\in\Theta_{p}$ into the same positional good level.
-The proof is in Appendix B.
+If $J(\theta)$ is not monotonic, define $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$ and let $\operatorname{conv}\tilde{J}$ denote the convex hull of $\tilde{J}$.
+Then, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$.
 Remark 3.
-Proposition 2 still holds under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1/2]$.
-To see this, when $\gamma<1/2$ (e.g., Immorlica et al.
-(2015) where $\gamma=0$), pooling is even less likely to be optimal as it reduces expected total status.
-161616 Formally, fix a positional good allocation $\chi$, the induced interim status $\tilde{s}(\theta)$ when $\gamma<1/2$ and $s(\theta)$ under the original specification ($\gamma=1/2$) satisfy $\tilde{s}(\theta)\leq s(\theta)$ for all $\theta\in[\theta_{0},\bar{\theta}]$, with equality if and only if $\tilde{s}(x)=F(x)$ for all $x\in[\theta_{0},\theta]$.
-Remark 6.
-Under the signaling specification $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$ (see Remark 4), Proposition 2 also holds (with a different optimal cutoff $\theta^{*}_{0}$) because the only difference is that $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$ instead.
-Remark 7.
-When restricted to deterministic mechanisms and under the specification $\tilde{S}({q},G)=\phi(S({q},G))$, where $\phi$ is an increasing and convex function, Proposition 2 still holds (with a different $\theta^{*}_{0}$) because Fan-Lorentz Theorem still applies (see details in the proof).
+Proposition 2 still holds under the more general specification $S(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1/2]$.
+To see this, when $\gamma<1/2$ (e.g., Immorlica, Stoddard and Syrgkanis (2015) where $\gamma=0$), pooling is even less likely to be optimal as it reduces expected total status.
 The revenue-maximizing mechanism can be implemented by an all-pay auction with a reserve price, where the higher the buyer pays (above the reserve price), the higher-level positional good he receives.
 In applications such as education or social status, it can be interpreted as an all-pay contest in which those whose effort does not meet the reserve price are punished with exclusion (expulsion from the organization).
-If the virtual valuation $J(\theta)$ is decreasing for some types, the revenue-maximizing mechanism pools these types into the same positional good level.
-In general, the revenue-maximizing mechanism is derived by applying the ironing technique to obtain the convex hull of $\tilde{J}(\theta)\equiv\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$.
+In general, the revenue-maximizing mechanism is derived by applying the ironing technique to obtain the convex hull of $\tilde{J}(\theta)$.
+If the virtual valuation $J(\theta)$ is decreasing for some types, the revenue-maximizing mechanism pools them into the same positional good level along with some adjacent types.
 The proposition has several implications.
 First, under Myerson’s regularity condition, it is optimal to sell as many levels of positional goods as possible.
-This is because as the number of position levels increases, the seller can induce a finer partition of buyers and thus extract a higher revenue.
-171717 For any two increasing allocations $s,\hat{s}\in\operatorname{\text{MPS}}(F)$ on $[\theta_{0},\bar{\theta}]$, if $s$ is finer than $\hat{s}$, then $\hat{s}\in\operatorname{\text{MPS}}(s)$ (the converse is not true).
+This is because, as the number of position levels increases, the seller can induce a finer partition of buyers and thus extract a higher revenue.
+[^17] For any two increasing allocations $s,\hat{s}\in\operatorname{\text{MPS}}(F)$ on $\theta_{0},\bar{\theta}]$, if $s$ is finer than $\hat{s}$, then $\hat{s}\in\operatorname{\text{MPS}}(s)$ (the converse is not true).
 Thus, I also say $s$ is finer than $\hat{s}$ in terms of majorization if $\hat{s}\in\operatorname{\text{MPS}}(s)$.
@@
 Second, compared to selling ordinary goods with intrinsic value $v(\theta)$ only, attaching the status value benefits the seller.
-This explains why luxury companies invest heavily in marketing to create the status value.
+This might explain why luxury companies invest heavily in marketing to create the status value.
 Third, the seller also benefits from excluding low types.
 For uniform distribution on $[0,1]$ and when $v(\theta)=0$, exclusion increases the revenue by 23.5%.
-Thus, exclusion is important for the seller (organization designer) to extract more revenue (effort).
+Thus, exclusion is important for the seller to extract more revenue.
 In applications where exclusion is impossible, full separation remains optimal under the regularity condition (See Proposition A.1).
@@
 Furthermore, when the intrinsic value $v(0)$ is large enough and there are many low types ($f(0)$ is large), the seller cannot benefit from exclusion.
-Approximation.
+3.1.1 Approximation
 In some applications, the number of positional good levels the seller can offer is subject to regulation, practical constraints, or other costs.
-The following proposition shows that the seller can still obtain a substantial portion of the maximum revenue by offering only one level of positional good.
+The following proposition shows that the seller can still obtain a substantial portion of the maximum revenue by offering a single good.
 Proposition 3.
-The seller can obtain at least half the maximum revenue by selling one level of positional good.
-{proofsketch}
+The seller can obtain at least half the maximum revenue by selling a single good.
+Proof sketch.
 Consider an auxiliary screening problem of selling indivisible ordinary goods (e.g., apples), where a posted price $p^{*}=\operatorname*{arg\,max}_{p}p(1-F(p))$ is optimal.
-I show that (i) by selling one level of positional good, the seller can obtain at least $\frac{1}{2}\max p(1-F(p))$ because those who buy the good get at least a status level of $s=1/2$; and (ii) the maximum revenue from selling positional goods is strictly below $\max p(1-F(p))$ because of positional externalities.
-(See the complete proof in Appendix B.)
+I show that (i) by selling a single good, the seller can obtain at least $\frac{1}{2}\max p(1-F(p))$ because those who buy the good get at least a status level of $s=1/2$; and (ii) the maximum revenue from selling positional goods is strictly below $\max p(1-F(p))$ because of positional externalities.
+(See the complete proof in Appendix B.) ∎
 Remark 5.
@@
 In the screening problem, selling an item to one consumer has no externalities on others.
-By contrast, negative externalities are present when selling positional goods—the seller cannot offer the highest status to multiple buyers at face value.
-Thus, the seller of positional goods is subject to an additional constraint, resulting in strictly less revenue.
+By contrast, when selling positional goods, because consumers care about their relative positions, negative externalities are present—the seller cannot offer the highest status to multiple buyers without reducing its value.
+Thus, the seller of positional goods is subject to an additional feasibility constraint, resulting in a strictly lower revenue.
 Graphical Illustration.
-Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$.
-Figure 2 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$.
-Assume the intrinsic value $v(\theta)=0$ for simplicity.181818 A positive intrinsic value $v(\theta)\geq 0$ will increase $\frac{R_{1}}{\max R}$ by increasing both $R_{1}$ and $\max R$.
-In Figure 2, (i) the blue area in the left panel represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$.
-(ii) The red area in the right panel represents the revenue from selling one level of positional good at price $F(\theta_{0}^{*})$—i.e., $\bar{s}(\theta)=\frac{1+F(\theta_{0}^{*})}{2}\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$.
-191919 To see (i) and (ii), the revenue under $s^{*}$ is $\int_{\theta_{0}^{*}}^{1}J(x)F(x)\mathop{}\!\mathrm{d}F(x)=-\int_{\tau^{*}_{0}}^{1}\tau\mathop{}\!\mathrm{d}R(\tau)=\tau^{*}_{0}R(\tau^{*}_{0})+\int_{\tau^{*}_{0}}^{1}R(\tau)\mathop{}\!\mathrm{d}\tau$.
-Analogously, the revenue under $\bar{s}$ is $\int_{\theta_{0}^{*}}^{1}J(x)\frac{1+F(\theta_{0}^{*})}{2}\mathop{}\!\mathrm{d}F(x)=\frac{1+\tau^{*}_{0}}{2}R(\tau^{*}_{0})$.
-Because the seller can optimize the price (and setting it at $F(\theta_{0}^{*})$ is feasible), the revenue from selling one level of positional good is at least the red area.
-(iii) The entire box in either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem.
-It is straightforward to see that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
+Figure 2 plots the revenue curve $R(\tau)=(1-\tau)F^{-1}(\tau)$ in the quantile space $\tau=F(\theta)$.
+Assume the intrinsic value $v(\theta)=0$ for simplicity.
+[^18] A positive intrinsic value $v(\theta)\geq 0$ will increase the ratio $\frac{R_{1}}{\max R}$, as it increases $R_{1}$ by at least the same amount as $\max R$.
+Then, the blue area in Figure 2 represents the revenue from the revenue-maximizing $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$.
+To see this, note that $R^{\prime}(\tau)=-J(\tau)$, so using a change of variable $\tau=F(\theta)$ (and $\tau_{0}^{*}=F(\theta_{0}^{*})$) and integration by parts, the revenue under $s^{*}(\theta)$ is
+$\int_{\theta_{0}^{*}}^{1}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)=-\int_{\tau^{*}_{0}}^{1}\tau\mathop{}\!\mathrm{d}R(\tau)=\tau^{*}_{0}R(\tau^{*}_{0})+\int_{\tau^{*}_{0}}^{1}R(\tau)\mathop{}\!\mathrm{d}\tau,$
+which is the area of the blue region.
 Figure 2: Revenue curve $R(\tau)$ for the uniform distribution
+Moreover, the red trapezoid in the right panel represents the revenue from selling a single good at price $F(\theta_{0}^{*})$—i.e., $\bar{s}(\theta)=\frac{1+F(\theta_{0}^{*})}{2}\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$.
+This is because the revenue under $\bar{s}$ is
+$\int_{\theta_{0}^{*}}^{1}J(\theta)\frac{1+F(\theta_{0}^{*})}{2}\mathop{}\!\mathrm{d}F(\theta)=\frac{1+\tau^{*}_{0}}{2}R(\tau^{*}_{0}),$
+which is the area of the red trapezoid in the right panel.
+Because the seller can optimize the price (and setting it at $F(\theta_{0}^{*})$ is feasible), the revenue from selling a single good is at least the red area.
+Finally, the entire box, which has an area of $R(\tau^{*}_{0})$, represents the maximum revenue from selling nonpositional goods in the auxiliary problem.
+It is straightforward to see that the red area is larger than half of the entire box and thus strictly larger than half of the blue area.
 The approximation is very close for many common distributions, as shown in the following examples (with $v(\theta)=0$).
 Example (Exponential distribution).
-If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, selling one level of positional good can obtain $91.9\%$ of the maximum revenue.
+If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, selling a single good can obtain $91.9\%$ of the maximum revenue.
 Example (Uniform distribution).
-If $F(\theta)=\theta$ on $[0,1]$, selling one level of positional good can obtain $92.4\%$ of the maximum revenue.
+If $F(\theta)=\theta$ on $[0,1]$, selling a single good can obtain $92.4\%$ of the maximum revenue.
 Example (Power distribution).
-If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, selling one level of positional good can obtain
+If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, selling a single good can obtain
 $\frac{R_{1}}{\max R}=\frac{(1+\beta)\left(\frac{1}{1+2\beta}\right)^{\frac{1}{2\beta}}}{\beta+(1+\beta)^{-1-\frac{1}{\beta}}}>91.4\%$
@@
 In particular, do full separation and exclusion also benefit consumers?
-The consumer surplus (or agent welfare) is given by
+The consumer surplus (or agents’ welfare) is given by
 $W=\int_{\theta_{0}}^{\bar{\theta}}U(\theta)\mathop{}\!\mathrm{d}F(\theta)=\int_{\theta_{0}}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)(s(\theta)+v^{\prime}(\theta))+U(\theta_{0})\mathop{}\!\mathrm{d}F(\theta)$ (11)
@@
   1. (i)
-    Fixing exclusion $\theta_{0}$, 202020 The exclusion level is fixed when the price of the lowest-level good is fixed.
+    Fixing exclusion $\theta_{0}$,  [^19] The exclusion level is fixed when the price of the lowest-level good is fixed.
 In particular, when there is a free regular service, there is no exclusion (see Appendix A.2).
@@
 This is because when the seller offers more levels of positional goods, consumer participation also increases, which in turn increases the consumer surplus under IFR.
-It is also worth noting that under DFR, if the distribution still satisfies Myerson’s regularity, the seller’s revenue and the consumer surplus are aligned—both are higher when the seller offers more levels of positional goods.
-Consumer Surplus Maximization.
-Relatedly, I study the optimal mechanism that maximizes consumer surplus. After rewriting the consumer surplus as
+It is also worth noting that if the distribution satisfies Myerson’s regularity but has DFR (i.e., a heavy tail), the seller’s revenue and the consumer surplus are aligned—both are higher when the seller offers more levels of positional goods.
+Intuitively, offering more levels of positional goods benefits high types and hurts low types, so it benefits agents in aggregate when there are relatively more high types.
+3.2.1 Consumer Surplus Maximization
+Consider next the optimal mechanism that maximizes consumer surplus. After rewriting the consumer surplus as
 $W=\int_{\theta_{0}}^{\bar{\theta}}U(\theta)\mathop{}\!\mathrm{d}F(\theta)=\int_{\theta_{0}}^{\bar{\theta}}(\theta s(\theta)-p(\theta)+v(\theta))\mathop{}\!\mathrm{d}F(\theta),$ (12)
 it is straightforward to show that when negative transfers are allowed and subject to budget balance ($\operatorname{\mathbf{E}}[p(\theta)]=0$), the consumer surplus is maximized by full separation without exclusion.
-Intuitively, this is because assortative matching is efficient, and transfers can redistribute efficiency gains among consumers.212121 See also Gershkov and Schweinzer (2010).
+Intuitively, this is because assortative matching is efficient, and transfers can redistribute efficiency gains among consumers and induce truth-telling.
+[^20] See also Gershkov and Schweinzer (2010).
 I thank Alex Gershkov for pointing this out.
@@
 Nonnegative Prices.
-Negative prices are implausible in some applications.
+The above result relies on the assumption that negative transfers are allowed, which is implausible in some applications.
 In particular, in the applications to organizational hierarchies and education, the price (i.e., effort) must be nonnegative.
@@
 Under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$, the first (second) part still holds if $\gamma\in[1/2,1]$ ($\gamma\in[0,1/2]$).
-In particular, full separation is always optimal if $\gamma=0$ (Immorlica et al., 2015), while total pooling is always optimal if $\gamma=1$ (Frank, 1985b).
+In particular, full separation is always optimal if $\gamma=0$ (Immorlica, Stoddard and Syrgkanis, 2015), while total pooling is always optimal if $\gamma=1$ (Frank, 1985b).
 Remark 7.
@@
 The result shows that the marginal loss of the excluded consumers exceeds the inframarginal gain of participants, so exclusion is detrimental to the consumer.
-On the other hand, under DFR, exclusion has no benefits to participants because the optimal mechanism fully separate participants, so no exclusion is optimal.
-The result implies that under IFR, the consumer-optimal regulation of positional goods without using subsidies is to restrict the seller to offer one level of positional good and serve all consumers (e.g., by providing a free regular boarding service).
+On the other hand, under DFR, exclusion has no benefits to participants because the optimal mechanism fully separates participants, so no exclusion is optimal.
+The result implies that under IFR, the consumer-optimal regulation of positional goods without using subsidies is to restrict the seller to offer a single good and serve all consumers (e.g., by providing a free regular boarding service).
 The results also have several implications in organizational hierarchies and education, where the effort (price) has to be positive.
-On the one hand, if the type distribution satisfies the IFR property (i.e, has a thin tail), which is satisfied by many common distributions, completely eliminating status competitions (e.g., educational disarmament) is optimal for agent welfare.
-In other words, agents pay nothing and are pooled at the same status.
+On the one hand, if the type distribution satisfies the IFR property (i.e, has a thin tail), completely eliminating status competitions (e.g., educational disarmament) is optimal for agents’ welfare.
 In the education application, this implies complete educational disarmament—for example, random assignment in allocating slots at sought-after schools.
-On the other hand, if there are a few very high types (i.e., a heavy tail), such as the Pareto distribution, intense competition or meritocracy is optimal even from the agent’s perspective.
+On the other hand, if there are a few very high types (i.e., a heavy tail), such as the Pareto distribution, meritocracy is optimal for agents’ welfare.
+Intuitively, low types benefit from soft competition, while high types prefer intense competition; hence, on average, more (less) intense competition is optimal for agents’ welfare when the population contains relatively more high (low) types.
 3.3 Social Planner
-Assume that the social planner maximizes a weighted sum of the seller’s revenue and consumer surplus, so social welfare is $S=\int_{0}^{\bar{\theta}}\lambda p(\theta)+U(\theta)\mathop{}\!\mathrm{d}F(\theta)$, where $\lambda\geq 0$.
+Assume that the social planner maximizes a weighted sum of the seller’s revenue and consumer surplus, so social welfare is $W_{S}=\int_{0}^{\bar{\theta}}\lambda p(\theta)+U(\theta)\mathop{}\!\mathrm{d}F(\theta)$, where $\lambda\geq 0$.
 In the monopolist provision problem where $p$ is transferred to the seller, $\lambda$ captures the weight of the seller’s revenue relative to the consumer surplus.
-When the price is allowed to be negative, I assume $\lambda\geq 1$, because otherwise the social planner would set $p(\theta)=-\infty$ to transfer all money from the seller to consumers.
+When the price is allowed to be negative, I assume $\lambda\geq 1$, because otherwise the social planner would transfer all money from the seller to consumers.
 In the education or the organization application, where $p$ is the agent’s nonnegative effort (so utilities are nontransferable), $\lambda$ captures the extent to which effort is socially productive.
-Agent effort is completely unproductive when $\lambda=0$ and productive when $\lambda=1$.
+agents’ effort is completely unproductive when $\lambda=0$ and productive when $\lambda=1$.
 When $\lambda>1$, the social benefit of effort is higher than the private benefit.
-Denote $J_{\lambda}(\theta)=\lambda\theta-(\lambda-1)\frac{1-F(\theta)}{f(\theta)}$. By previous results, social welfare is
-$S=\int_{\theta_{0}}^{\bar{\theta}}J_{\lambda}(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{0}}^{\bar{\theta}}\left(\lambda v(\theta)-(\lambda v^{\prime}(\theta)-1)\frac{1-F(\theta)}{f(\theta)}-(\lambda-1)U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta).$ (13)
+Define $J_{\lambda}(\theta)=\lambda\theta-(\lambda-1)\frac{1-F(\theta)}{f(\theta)}$. The expected social welfare is
+$W_{S}=\int_{\theta_{0}}^{\bar{\theta}}J_{\lambda}(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{0}}^{\bar{\theta}}\left(\lambda v(\theta)-(\lambda v^{\prime}(\theta)-1)\frac{1-F(\theta)}{f(\theta)}-(\lambda-1)U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta).$ (13)
 Using the same argument as before, we have the following result.
 Proposition 6.
-The social welfare-maximizing mechanism fully separates the participants and potentially exclude some types
+The social welfare-maximizing mechanism fully separates the participants and potentially excludes some types
   1. (i)
@@
 4.1 Intrinsic Quality
-In this subsection, I consider the case where the positional good has an intrinsic quality ${q}$, produced at a cost $c({q})$, which determined its intrinsic value $v({q},\theta)=\theta q$ (à la Mussa and Rosen (1978)).
-Assume $c({q})$ is strictly increasing and strictly convex.
+I have maintained the assumption that the positional good has no intrinsic value to abstract from quality differentiation.
+However, in many applications, a higher-level positional good (e.g., a larger car or higher boarding class) also comes with a higher intrinsic value.
+Now assume the positional good q{q} has an intrinsic quality produced at a cost c(q)c({q}), which determines its intrinsic value $v({q},\theta)=\theta q$, à la Mussa and Rosen (1978).
+Assume c(q)c({q}) is strictly increasing and strictly convex.
 Denote by $Q(\theta)=\operatorname{\mathbf{E}}_{\omega}[\chi(\theta,\omega)]$ the interim allocation.
@@
 Proposition 7.
-Assume $J(\theta)$ is increasing. The revenue-maximizing mechanism excludes types below $\theta^{*}_{0}=J^{-1}(0)$ and offers $\chi^{*}(\theta)=c^{\prime-1}(J(\theta))$ to $\theta\geq\theta^{*}_{0}$.
+Assume $J(\theta)$ is increasing. The revenue-maximizing mechanism excludes types below $\theta^{*}_{0}=J^{-1}(0)$ and offers $Q^{*}(\theta)=c^{\prime-1}(J(\theta))$ to $\theta\geq\theta^{*}_{0}$.
 The proposition implies that the optimal mechanism in revenue maximization is robust to the introduction of intrinsic quality.
@@
 This explains the amount of money luxury companies spent on marketing (to create status value) instead of research and development (to reduce the cost).
+4.2 Convex or Concave Status
+When positional externalities arise from interpersonal comparison, a concave or convex specification of status may be plausible. Formally, assume the status function is
+$\tilde{S}({q},G)=\phi(S({q},G))=\phi\left(\frac{G^{-}({q})+G({q})}{2}\right),$ (15)
+where $\phi(\cdot)$ is a strictly increasing function.
+Under this specification, an increasing interim status $\tilde{s}$ is feasible if and only if $\tilde{s}\in\operatorname{conv}\{\sigma=\phi\circ s\mid s\in\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)\}$.
+Because the maximization problems remain linear in $\tilde{s}$, the optimal $\tilde{s}^{*}$ is an extreme point of the feasible set, which satisfies $\tilde{s}^{*}=\phi\circ s^{*}$ for some $s^{*}\in\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)$.
+Hence, the following proposition shows that the previous results are robust to concave or convex status, although they may require stronger regularity conditions.
+Proposition 8.
+(i) If $\phi$ is strictly increasing and convex, Propositions 2, 4, 5, and 6 still hold (with a different exclusion level) after replacing the IFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is decreasing) with
+$\frac{1-F(\theta)}{f(\theta)}\phi^{\prime}(F(\theta))\mbox{ is decreasing.}$
+(ii) If $\phi$ is strictly increasing and concave, Propositions 2, 4, 5, and 6 still hold (with a different exclusion level) after replacing
+  • Myerson’s regularity (i.e., $J(\theta)$ is increasing) with
+    $J(\theta)\phi^{\prime}(F(\theta))\mbox{ is increasing;}$
+  • the DFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is increasing) with
+    $\frac{1-F(\theta)}{f(\theta)}\phi^{\prime}(F(\theta))\mbox{ is increasing;}$
+  • $J_{\lambda}(\theta)$ is increasing with
+    $J_{\lambda}(\theta)\phi^{\prime}(F(\theta))\mbox{ is increasing.}$
+In cases where conditions are unchanged (e.g., in revenue maximization if $\phi$ is convex), the result follows from the Fan-Lorentz Theorem.
+When conditions are stronger, the proof uses the optimal control method.
 4.3 Deliberately Lowering the Status
 In some applications, the principal can deliberately lower the status.
-In the priority service example (Example 2.1), the seller can prolong the waiting time $t=1-s$ by pausing service or procrastinating.
-Thus, the seller can create negative status $s<0$ or excessive waiting time $t>1$.
-222222 For example, in the U.S., OPT processing time is excessively long for the regular service but takes only 30 days for the premium service.
-In these cases, the range of status allocation $s$ becomes $[-M,1]$, where $M$ is a sufficiently large bound on the negative status (or the maximum excessive waiting time).
+In the priority service example (Example 1), the seller can prolong the waiting time t=1st=1-s by pausing service or procrastinating.
+Thus, the seller can create excessive waiting time $t>1$ or essentially negative status $s<0$.
+In these cases, the range of status allocation $s$ becomes [M,1][-M,1], where $M>0$ is a sufficiently large bound on the negative status (or the maximum excessive waiting time).
 Incentive compatibility still requires $s$ to be increasing, and feasibility requires $s\leq\hat{s}$ for some $\hat{s}\in\operatorname{\text{MPS}}_{w}(F)$.
-By lowering the status, the seller can offer negative status $s<0$ to low types and reduces their payoffs.
+By lowering the status, the seller can offer negative status $s<0$ to low types and reduce their payoffs.
 Therefore, negative status (or excessive waiting time) serves as a substitute for exclusion.
@@
 Under IFR, allowing the seller to lower the status increases the seller’s revenue and decreases the consumer surplus.
-232323 It strictly increases the seller’s revenue if $v$ is nonlinear (i.e., strictly concave or $v(0)>0$) and strictly decreases the consumer surplus if $v$ is nonlinear and $f(0)<\infty$.
+[^21] It strictly increases the seller’s revenue if $v$ is nonlinear (i.e., strictly concave or $v(0)>0$) and strictly decreases the consumer surplus if $v$ is nonlinear and $f(0)<\infty$.
 Remark 8.
-The allocation $s^{*}(\theta)=-v^{\prime}(\theta)$ is increasing because $v^{\prime\prime}(\theta)\leq 0$ and feasible because it can be induced by randomization.
-Example 4.1 (Excessive Waiting Time).
-Consider the payoff $u(p,s,\theta)=\theta s-p+v(\theta)$ ($=\tilde{v}(\theta)-\theta t-p$), where $t=1-s$ is the waiting time.
+The allocation $s^{*}(\theta)=-v^{\prime}(\theta)$ is increasing because $v^{\prime\prime}(\theta)\leq 0$ and feasible because it can be induced by randomization between pooling and separation (after lowering the status by a constant).
+Example 3 (Excessive Waiting Time).
+Consider the payoff $u(p,s,\theta)=\theta s-p+v(\theta)$ ($\tilde{v}(\theta)-\theta t-p$), where t=1st=1-s is the waiting time.
 Assume $\theta\sim\text{Unif }[0,1]$ and that $v(\theta)=v_{0}+\alpha\theta$, where $v_{0},\alpha>0$.
@@
 Figure 3: Excessive Waiting Time
-The insight that negative status serves as a substitute for exclusion also applies to the case where exclusion is impossible as the lowest-level good or service is free (see Appendix A).
+The insight that negative status serves as a substitute for exclusion also applies to the case where exclusion is impossible, as the lowest-level good or service is free (see Appendix A).
 In this case, negative status or excessive waiting time diminishes the value of the free service and effectively plays the role of exclusion, thereby increasing the seller’s revenue.
 In particular, by offering a priority service and a free regular service, the seller extracts more revenue as she increases the waiting time for the free regular service.
+4.4 Signaling
+Consider an alternative specification where the status arises from signaling concerns.
+For example, consumers buy conspicuous goods $q$ to obtain social status with a value of $\operatorname{\mathbf{E}}[\theta\mid q]$ (Rayo, 2013).
+Under this specification, the status value induced by a positional good allocation $\chi(\theta)$ is given by $S({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$.
+Moreover, even if the consumer does not buy the good (i.e., $q=0$), he can still obtain some status value $\underline{s}=\operatorname{\mathbf{E}}[\theta\mid q=0]$, instead of zero in the main specification.
+Therefore, exclusion is impossible because the consumer can get the lowest status $\underline{s}$ for free.
+Analogous to Theorem 1 (without exclusion), an increasing interim status $\tilde{s}$ is feasible if and only if $s\in\operatorname{\text{MPS}}(\theta)$ in the quantile space.
+Therefore, the main results continue to hold except that exclusion is impossible (see also Appendix A).
+In particular, the revenue-maximizing mechanism fully separates all types if and only if $J(\theta)$ is increasing, while selling more levels reduces the consumer surplus if $F$ satisfies the IFR property.
 4.5 Status Game As a Suffering
-In this subsection, I consider the case where $v^{\prime}(\theta)\leq-1$, so the status game is a “suffering” in the sense that participants with higher types suffer more and receive lower payoffs.
-In the queuing example (Example 2.1), $v^{\prime}(\theta)\leq-1$ holds if higher types derive lower utility from the service after accounting for waiting costs—for example, if the service has the same value $v_{0}$ to everyone: $u(p,s,\theta)=v_{0}-\theta t-p$, where $t=1-s$ is the waiting time (and thus $v(\theta)=v_{0}-\theta$).
-Because higher types face higher costs or disutility, agents have incentives to overreport their types to get lower waiting time.
+Now I consider a case where $v^{\prime}(\theta)\leq-1$, so that the status game is a “suffering” in the sense that higher types suffer more and receive lower payoffs when they participate.
+In the queuing example (Example 1), if the good they are waiting for has homogenous values across agents, higher (cost) types will derive lower utility from it after accounting for waiting costs.
+[^22] In Example 1, if the good has the same value $v_{0}$ to everyone, then the payoff $u(p,s,\theta)=v_{0}-\theta t-p=(v_{0}-\theta)+\theta s-p$ (because t=1st=1-s is the waiting time) satisfies $v^{\prime}(\theta)=-1$.
+In general, because higher types face higher costs or disutility, agents have incentives to overreport their types to get lower waiting time.
 Formally, because $U^{\prime}(\theta)=s(\theta)+v^{\prime}(\theta)\leq 0$ for all participants when $v^{\prime}(\theta)\leq-1$, higher types have lower payoffs, thereby creating incentives for overreporting, in contrast to underreporting in the main specification.
@@
   1. (i)
-    Assume $L(\theta)=\theta+F(\theta)/f(\theta)$ is increasing. Then,
+    Assume $L(\theta)=\theta+\frac{F(\theta)}{f(\theta)}$ is increasing. Then,
     1. (a)
@@
     3. (c)
-      If $-v^{\prime}(\theta)/v(\theta)\leq f(\theta)/F(\theta)$ for all $\theta\in[0,\bar{\theta}]$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=\bar{\theta}$).
+      If $\frac{v^{\prime}(\theta)}{v(\theta)}\leq\frac{f(\theta)}{F(\theta)}$ for all $\theta\in[0,\bar{\theta}]$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=\bar{\theta}$).
   2. (ii)
-    The seller can obtain at least half the maximum revenue by selling one level of positional good.
+    The seller can obtain at least half the maximum revenue by selling a single good.
   3. (iii)
@@
     Deliberately lowering the status cannot increase the seller’s revenue.
-Parts (i)–(iii) are qualitatively similar to those in the main specification, except that the revenue-maximizing mechanism excludes high types instead of low types.
-For part (iv), when the seller assign negative status, she has to assign it to low types first because $s$ is increasing (by incentive compatibility).
-Hence, because $U^{\prime}(\theta)\leq 0$, negative status reduces the payoffs of all participants, and therefore cannot increase the seller’s revenue.
+Parts (i)–(iii) are qualitatively similar to those in the main specification, with the integrand $J(\theta)$ replaced by $L(\theta)=\theta+\frac{F(\theta)}{f(\theta)}$, and that the revenue-maximizing mechanism excludes high types instead of low types (because $U^{\prime}(\theta)\leq 0$).
+For part (iv), when the seller assigns a negative status, she must assign it to low types first because $s$ is increasing.
+However, since agents now have incentives to overreport (because $U^{\prime}(\theta)\leq 0$), assigning negative status to low types does not mitigate agents’ incentives to overreport and therefore cannot increase the seller’s revenue.
 Countervailing incentives.
@@
 5 Conclusion
-In this paper, I study the allocation of positional goods using a mechanism approach.
-Many goods are positional in the sense that allocating an item to one agent has externalities on others: moving up the position of one consumer would push down that of another, and the value of the good depreciates as more people have it.
-The model has dual implications in the monopolist provision of positional goods and the design of status categories within organizations.
+In this paper, I study the allocation of positional goods using a mechanism design approach.
+Many goods are positional because consumers care about their relative positions in consumption.
+Thus, allocating an item to one agent has externalities on others: moving up the position of one consumer would push down that of another, and the value of the good depreciates as more people have it.
+Because of these externalities, positional goods remain scarce and competitive despite economic growth.
+My model has implications for both the monopolist provision of positional goods and the design of status categories within organizations.
+In terms of the monopolist problem, the seller can often extract more revenue by offering many levels of positional goods through their status values.
+At the extreme, an all-pay auction with a reserve price, where the more a buyer pays, the higher position they receive, can maximize the seller’s revenue.
+By contrast, offering more levels of positional goods is often detrimental to consumers.
+Thus, it is usually beneficial for consumers to restrict the number of positional good levels the seller can offer, and the consumer-optimal regulation without using subsidies is to restrict the seller to offering a single good and serving everyone.
+Nevertheless, in cases where a few buyers have very high valuations (e.g., Pareto distribution), both the seller’s revenue and consumer surplus are higher when more status levels are offered—their objectives are aligned.
 The results also shed light on the trade-offs involved in educational disarmament.
-In terms of the monopolist problem, the seller can often extract more revenue by offering many levels of positional goods through their status values.
-At the extreme, an all-pay auction with reserve prices, where the more buyers pay, the higher positions they receive, can maximize the seller’s revenue.
-On the other hand, offering more levels of positional goods is often detrimental to consumers.
-Thus, it is usually beneficial for consumers to restrict the number of position levels the seller can offer, and the consumer-optimal regulation without using subsidies is to restrict the seller to only one levels of positional goods and serving everyone.
-Nevertheless, in cases where a few buyers have very high valuations (e.g., Pareto distribution), both the seller’s revenue and the consumer surplus are higher when more status levels are offered—their objectives are aligned.
-In terms of the organizational design problem, the results suggest that the management practice of stack ranking and rat races in education maximizes agents’ effort, consistent with the common wisdom, if the ability distribution satisfies Myerson’s regularity.
-The exclusion of low types can increase revenue or effort.
-On the contrary, from the agent’s perspective, less intense status competition is beneficial if the ability distribution satisfies IFR (or has a thin tail).
-This explains the call for abolishing stack ranking and educational disarmament, such as lotteries for school admissions.
-Nevertheless, when a few agents have very high abilities (e.g., Pareto distribution), rat races can be better for agent welfare.
-Thus, the objectives to maximize agent effort and welfare can be aligned, which challenges the common wisdom that rat races are always detrimental to agents.
-In either case, exclusion always harms agent welfare, even though it can increase the status of participants.
+The results suggest that rat races in education maximize agents’ effort, consistent with the common wisdom, if the ability distribution satisfies Myerson’s regularity.
+By contrast, from the agent’s perspective, reduced status competition is beneficial if the ability distribution satisfies IFR (i.e., has a thin tail).
+This explains calls for educational disarmament, such as lotteries for school admissions.
+Nevertheless, when a few agents have very high abilities (e.g., Pareto distribution), educational arms races can be better for agents’ welfare.
+Intuitively, high-ability (low-ability) agents benefit from more (less) competition, so meritocracy raises (reduces) agents’ welfare on average when there are relatively more high-ability (low-ability) agents.
+Thus, the objectives of maximizing agents’ effort and welfare can be aligned, which challenges the common wisdom that rat races are always detrimental to agents.
+In either case, exclusion always harms agents’ welfare, even if it raises the status of participants.
+Similarly, in organizations, the results suggest that the management practice of stack ranking maximizes agents’ effort under Myerson’s regularity but harms agents’ welfare under IFR.
 Moreover, when learning effort is productive, education arms races can increase social welfare.
@@
 This can be interpreted as offering a different track of education that leads to different status levels (e.g., vocational versus academic education) or having another product that has a separate queue.
-In either case, consumers of one set of positional goods derive utility independently of those of the other, either because it is psychologically incomparable or physically separate.
-If the other seller offers the same menu as the original seller, an immediate result is that each seller will attract half of the buyers, and that competition can increase consumer surplus.242424 Di Stasio et al.
-(2016) find that education is less positional in countries with more developed vocational education systems (e.g., Germany).
-However, when sellers compete in mechanisms strategically, the problem remains an open question.
+In either case, consumers of one set of positional goods derive utility independently of those of the other, either because they are psychologically incomparable or physically separate.
+If the other seller offers the same menu, each seller will attract half of the buyers, and that competition can increase consumer surplus.
+[^23] Di Stasio, Bol and Van de Werfhorst (2016) find that education is less positional in countries with more developed vocational education systems (e.g., Germany).
+However, when sellers compete in mechanisms strategically, the problem remains open.
 Appendix A Optimal Mechanisms without Exclusion
 In this section, I assume exclusion is impossible and that the lowest-level positional good is free.
-For example, in Moldovanu, Sela, and Shi (2007), agents cannot be excluded and will at least get the lowest status in the organization.
+[^24] This section subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.” For example, in Benny Moldovanu, Aner Sela and Xianwen Shi (2007), agents cannot be excluded and can get at least the lowest status in the organization.
 In the benchmark model of Gershkov and Winter (2023), buyers can use the regular (non-priority) service for free.
@@
 The (IR) constraint, which is equivalent to $\tilde{U}(\theta)\equiv U(\theta)-v(\theta)-\theta\underline{s}\geq 0$, can be reduced to $U(0)-v(0)\geq 0$ because $U^{\prime}(\theta)-v^{\prime}(\theta)-\underline{s}=s(\theta)-\underline{s}\geq 0$.
-The revenue is given by
+By standard arguments, the expected revenue is given by
 $R=\int_{0}^{\bar{\theta}}p(\theta)\mathop{}\!\mathrm{d}F(\theta)$ $=\int_{0}^{\bar{\theta}}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)s(\theta)\mathop{}\!\mathrm{d}F(\theta)-(U(0)-v(0)).$ (A.6)
-By (IR), it is optimal to set $U(0)=v(0)$. Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. The revenue maximization problem is equivalent to
+By (IR), it is optimal to set $U(0)=v(0)$. The revenue maximization problem is equivalent to
 $\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta).$ (A.7)
@@
 Proposition A.1.
-Offering more positional good levels (strictly) increases the seller’s revenue if and only if $J(\theta)$ is (strictly) increasing.
+Offering more positional good levels always (strictly) increases the seller’s revenue if and only if $J(\theta)$ is (strictly) increasing.
 Thus, full separation maximizes revenue if and only if $J(\theta)$ is increasing.
+If $J(\theta)$ is not monotonic, define $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$ and let $\operatorname{conv}\tilde{J}$ denote the convex hull of $\tilde{J}$.
+Then, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$.
 Remark A.1.
-GW’s Proposition 8 provides a sufficient condition for the seller’s revenue to be strictly increasing in the number of priority classes: $F$ satisfies the IFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is decreasing).
+GW’s Proposition 8 provides a sufficient condition for the seller’s revenue to be strictly increasing in the number of priority classes: $F$ satisfies the IFR property.
 Remark A.2.
 Effort maximization in MSS’s model is the same as revenue maximization because they assume linear effort costs (and no exclusion).
-They provide a sufficient condition for full separation in Theorem 4: $F$ satisfies the IFR property.
+In Theorem 4, they also provide a sufficient condition for full separation : $F$ satisfies the IFR property.
 Infinitely many positional good levels can be implemented by an all-pay auction, in which the more money a consumer pays, the higher status he receives.
@@
 The revenue-maximizing mechanism always separates the highest types $\theta\in[\bar{\theta}-\varepsilon,\bar{\theta}]$ for some $\varepsilon>0$.
-{proofsketch}
-Because $J^{\prime}(\bar{\theta})=2>0$, $J(\theta)$ is increasing in a neighborhood of $\bar{\theta}$, so by Proposition 2 in KMS, types in this neighborhood are separated.
+Proof sketch.
+Because $J^{\prime}(\bar{\theta})=2>0$, $J(\theta)$ is increasing in a neighborhood of $\bar{\theta}$, so by Proposition 2 in KMS, types in this neighborhood are separated. ∎
 Remark A.3.
@@
 Proposition A.3.
-The seller can obtain at least half the maximum revenue by offering one positional good level in addition to a free low-level position.
+The seller can obtain at least half the maximum revenue by selling a single positional good in addition to a free low-level position.
 Proof.
@@
 $\theta(p)\frac{1+F(\theta(p))}{2}-p=\theta(p)\frac{F(\theta(p))}{2}\iff\theta(p)=2p.$
-By offering one positional good level in addition to a free low-level position, the seller’s maximum revenue is
+By offering one positional good in addition to a free low-level position, the seller’s maximum revenue is
 $R_{2}=\max_{p}p(1-F(2p))=\frac{1}{2}\max_{p}p(1-F(p)).$
@@
 Graphical Illustration.
-Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$.
-Figure 4 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$.
+Figure 4 plots the revenue curve $R(\tau)=(1-\tau)F^{-1}(\tau)$ in the quantile space $\tau=F(\theta)$.
 Assume $v(\theta)=0$ for simplicity.
-In Figure 4, (i) the area under the revenue curve (blue area in the left panel) represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)$, (ii) the red area in the right panel represents the revenue from having two levels of positional goods—a higher level offered at price $F(\theta^{*}_{0})$ in addition to a lower level offered for free, and (iii) the entire box of either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem.
+Similar to Section 3.1.1, (i) the area under the revenue curve (blue area in the left panel) represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)$, (ii) the red triangle in the right panel represents the revenue from having two levels of positional goods—a higher level offered at price $F(\theta^{*}_{0})$ in addition to a lower level offered for free, and (iii) the entire box of either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary problem.
 It is straightforward that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
@@
 Example (Exponential distribution).
-If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, offering one positional good level (in addition to a free low level) can obtain $73.6\%$ of the maximum revenue.
+If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, offering a single good above the free lowest level can obtain $73.6\%$ of the maximum revenue.
 Example (Uniform distribution).
-If $F(\theta)=\theta$ on $[0,1]$, offering one positional good level (in addition to a free low level) can obtain $75\%$ of the maximum revenue.
+If $F(\theta)=\theta$ on $[0,1]$, offering a single good above the free lowest level can obtain $75\%$ of the maximum revenue.
 Example (Power distribution).
-If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, offering one positional good level (in addition to a free low level) can obtain
+If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, offering a single good above the free lowest level can obtain
 $\frac{R_{2}}{\max R}=\begin{cases}\frac{(1+2\beta)(1+\beta)^{-1/\beta}}{2\beta}>72.1\%,&\text{ if }\beta\geq 1\\ \frac{\beta(1+2\beta)(1+\beta)^{-1/\beta}}{2\beta^{2}-(1+\beta)^{-\frac{\beta+1}{\beta(1-\beta)}}\left(2\beta^{2}+(1+\beta)^{-\frac{\beta}{1-\beta}}-(1+\beta)\right)}>\frac{3}{4},&\text{ if }0<\beta<1\end{cases}$
@@
 Therefore, the consumer surplus decreases (increases) as the seller offers more levels of positional goods if $F$ satisfies IFR (DFR).
+I focus on the case where the price (or effort) is nonnegative, which pin down p(0)=v(0)U(0)=0p(0)=v(0)-U(0)=0.
 Corollary A.4.1.
@@
 Remark A.5.
-If both $\theta-\frac{1-F(\theta)}{f(\theta)}$ and $\frac{1-F(\theta)}{f(\theta)}$ are increasing (e.g., exponential and Pareto distributions), a finer partition (in terms of majorization) can increase both the seller’s revenue and consumer surplus.
+If both $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$ and $\frac{1-F(\theta)}{f(\theta)}$ are increasing (e.g., exponential and Pareto distributions), a finer partition (in terms of majorization) can increase both the seller’s revenue and consumer surplus.
 Thus, full separation maximizes both the revenue and consumer surplus.
@@
 The consumer surplus is the highest when the seller offers one position level if and only if
-$\int_{0}^{\theta}\left(\frac{1-F(x)}{f(x)}-\operatorname{\mathbf{E}}[\theta]\right)\mathop{}\!\mathrm{d}F(x)\geq 0\mbox{ for all $\theta\in[0,\bar{\theta}]$.}$
+[…]
 Proof.
-Denote $I(\theta)=\int_{0}^{\theta}\frac{1-F(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)$.
-Then, the condition in the proposition is equivalent to $I(\theta)\geq I(\bar{\theta})F(\theta)=\operatorname{\mathbf{E}}[\theta]F(\theta)$ (graphically, $I(\theta)$ lies above the line connecting $I(0)=0$ and $I(\bar{\theta})=\operatorname{\mathbf{E}}[\theta]$ in the quantile space).
-Therefore, $I(\bar{\theta})F(\theta)$, which corresponds to total pooling, is the convex hull of $I$ (i.e., the largest convex function that lies below $I$).252525 A piecewise affine function consisting of other points on $I(\theta)$, which corresponds to offering more positional good levels, lies above $I(\bar{\theta})F(\theta)$ and leads to lower consumer surplus.
+Define $I(\theta)=\int_{0}^{\theta}\frac{1-F(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)$.
+Then, the condition in the proposition is equivalent to $I(\theta)\geq I(\bar{\theta})F(\theta)=\operatorname{\mathbf{E}}[\theta]F(\theta)$ (i.e., $I(\theta)$ lies above the line connecting $I(0)=0$ and $I(\bar{\theta})=\operatorname{\mathbf{E}}[\theta]$ in the quantile space).
+Therefore, $I(\bar{\theta})F(\theta)=\operatorname{conv}I$ (the largest convex function that lies below $I$).
 By Proposition 2 in KMS, this condition is necessary and sufficient for total pooling to be welfare-maximizing.
@@
 Moreover, fix a type $\theta\in[0,\theta]$, maximizing his utility $U(\theta)$ subject to constraints (A.2)–(A.5) is equivalent to
-$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)+U(0)-v(0).$ (A.9)
-Because $p(0)\geq 0$, it is optimal to set $p(0)=0$ and $U(0)=v(0)$.
+$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)$ (A.9)
 The following proposition provides a necessary and sufficient condition for total pooling to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
@@
 A sufficient condition is that $f(\theta)$ is increasing.
-{proofsketch}
+Proof sketch.
 $U(\theta)=\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)$. If $f$ is increasing, then $1/f$ is decreasing, so $s=1/2$ maximizes $U(\theta)$.
@@
 Similar to the proof of Proposition A.5, one can show that this condition is necessary and sufficient for $s(\theta)=1/2$ to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
+∎
 Remark A.8.
@@
 Appendix B Proofs
-B.1 Proof of Section 3
+B.1 Proof of Theorem 1
+Proof.
+($\implies$) follows from $\operatorname{\mathbf{E}}[s(\theta)\mid\theta\geq x]\leq\operatorname{\mathbf{E}}[F(\theta)\mid\theta\geq x]$ for all $x\in[0,\bar{\theta}]$ because switching from any $s(\theta)$ to assortative rule ($s(\theta)=F(\theta)$) takes the status of low type to high types.
+[…]
+$s(\theta)=\frac{1}{F(\bar{\theta}_{i})-F(\underline{\theta}_{i})}\int_{\underline{\theta}_{i}}^{\bar{\theta}_{i}}F(\theta)\mathop{}\!\mathrm{d}F(\theta)=\frac{F(\underline{\theta}_{i})+F(\bar{\theta}_{i})}{2};$
+and (iii) fully separating otherwise, where $s(\theta)=F(\theta)$.
+Hence, by Proposition 1 (Choquet’s Theorem) in KMS, there exists a probability measure $\lambda_{s}$ supported on the extreme points of $\operatorname{\text{MPS}}_{w}(F)$ such that $s=\operatorname{\mathbf{E}}[\tilde{s}\mid\tilde{s}\sim\lambda_{s}]$.
+Thus, $s$ can be induced by randomizing over allocations that each induces extreme points of $\operatorname{\text{MPS}}_{w}(F)$.
+∎
+B.2 Proofs of Section 3
+B.2.1 Proofs of Section 3.1
 Proof of Proposition 2.
-The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to seller’s revenue given $\theta_{0}$, that is,
+The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to
 $R(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)+v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)\right)\mathop{}\!\mathrm{d}F(\theta).$ (B.1)
@@
 $=\operatorname*{arg\,max}_{\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$
-Note that under the specification $\tilde{S}({q},G_{\chi})=\phi(S({q},G_{\chi}))$, where $\phi$ is an increasing and convex function, $J(\theta)s(\theta)$ in equation (B.1) is replaced by $J(\theta)\phi(s(\theta))$.
-Lemma B.1 still follows from Fan-Lorentz Theorem because $J(\theta)\phi(s)$ is convex in $s$ and supermodular in $(s,\theta)$ when $\phi$ is increasing and convex.
-On the other hand, if $J(\theta)$ is decreasing on $\Theta_{p}\subseteq[0,\bar{\theta}]$, Fan-Lorentz Theorem implies that the revenue-maximizing allocation pools all types $\theta\in\Theta_{p}$ at the same level.
+If $J(\theta)$ is not monotonic, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$.
 ∎
 Proof of Corollary 2.1.
-Denote $J_{v}(\theta)=v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)$.
+Define $J_{v}(\theta)=v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)$.
 Under IFR, $J_{v}(\theta)$ is increasing because $v(\theta),v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$.
-Denote
+Define
 $\psi(\theta)\equiv-R^{\prime}(\theta)/f(\theta)=J(\theta)F(\theta)+J_{v}(\theta).$ (B.2)
-Denote $\theta_{1}=J^{-1}(0)$. We have
+Define $\theta_{1}=J^{-1}(0)$. We have
 $\psi(\theta_{1})=J(\theta_{1})F(\theta_{1})+J_{v}(\theta_{1})=v(\theta_{1})-\theta_{1}v^{\prime}(\theta)\geq 0$ (B.3)
@@
 Proof of Proposition 3.
-Denoted by $p$ the price of the positional good. The cutoff type $\theta(p)$ indifferent between buying and not buying is given by
+Denote by $p$ the price of the positional good. The cutoff type $\theta(p)$ indifferent between buying and not buying is given by
 $\theta(p)\frac{1+F(\theta(p))}{2}-p+v(\theta(p))=0\implies p=\theta(p)\frac{1+F(\theta(p))}{2}+v(\theta(p)).$
-Because $v^{\prime}(\theta)\geq 0$, all types above $\theta(p)$ will buy the good. Thus, by selling one level of positional good, the seller’s maximum revenue is
+Because $v^{\prime}(\theta)\geq 0$, all types above $\theta(p)$ will buy the good. Thus, by selling a single good, the seller’s maximum revenue is
 $R_{1}=\max_{\theta}\left(\theta\frac{1+F(\theta)}{2}+v(\theta)\right)(1-F(\theta))\geq\frac{1}{2}\max_{\theta}(\theta+v(\theta))(1-F(\theta))$
@@
 Consider the auxiliary screening problem of selling an indivisible item of value to one buyer, in which a standard result implies a posted-price mechanism is optimal (see Börgers, 2015, Proposition 2.5).
-Formally, denote by $q$ the probability the item is allocated to $\theta$, and let =$\mathcal{M}=\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$.
+Formally, denote by $q$ the probability the item is allocated to $\theta$, and let $\mathcal{M}=\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$.
 The agent’s payoff is $\theta q-p$.
@@
 ∎
+B.2.2 Proofs of Section 3.2
 Proof of Proposition 4.
-(i) The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS.
+(i) The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to
+$W(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)(s(\theta)+v^{\prime}(\theta))+U(\theta_{0})\mathop{}\!\mathrm{d}F(\theta)$ (B.5)
 Lemma B.2.
@@
 If $s(\theta)$ is strictly increasing at $\theta_{0}$, then decreasing $\theta_{0}$ to $\theta_{0}-\varepsilon$ increases the consumer surplus of $\theta\in[\theta_{0}-\varepsilon,\theta_{0})$ and does not affect the consumer surplus of $\theta\in[\theta_{0},\bar{\theta}]$.
-If $s(\theta)=s_{0}$ is constant on $[\theta_{0},\theta_{1}]$ for some $\theta_{1}\in(\theta_{0},\bar{\theta}]$ (and $s(\theta)>s_{0}$ for $\theta>\theta_{1}$), then the consumer surplus is
+If $s(\theta)=s_{0}$ is constant on $\theta_{0},\theta_{1}]$ for some $\theta_{1}\in(\theta_{0},\bar{\theta}]$ (and $s(\theta)>s_{0}$ for $\theta>\theta_{1}$), then the consumer surplus is
 $W(\theta_{0})=\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+F(\theta_{1})/2+F(\theta_{0})/2\right)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{1}}^{\bar{\theta}}\frac{1-F(\theta)}{f(\theta)}\left(v^{\prime}(\theta)+s(\theta)\right)\mathop{}\!\mathrm{d}F(\theta),$
 with derivative given by
-[…]
-Denote
-$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta.$
-Note that $z(\theta_{1})-F(\theta_{1})-v^{\prime}(\theta_{1})=0$, and under IFR,
-$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
-Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}\leq\theta_{1}$.
+$\begin{aligned} W^{\prime}(\theta_{0})=-({1-F(\theta_{0})})(v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2)+\frac{f(\theta_{0})}{2}\int_{\theta_{0}}^{\theta_{1}}({1-F(\theta)})\mathop{}\!\mathrm{d}\theta.\end{aligned}$
+Define
+$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta$
+and $\psi(\theta)=z(\theta)-F(\theta)-v^{\prime}(\theta)$. Note that $\psi(\theta_{1})=0$, and under IFR,
+$\psi^{\prime}(\theta)=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
+Therefore, $\phi(\theta_{0})\geq 0$ for all $\theta_{0}\leq\theta_{1}$.
+It follows that $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$.
 Hence, decreasing $\theta_{0}$ increases the consumer surplus $W(\theta_{0})$.
@@
 with derivatives on $\theta_{0}>0$ given by
-[…]
-It is straightforward that $W^{\prime}(\bar{\theta})=0$ and $W^{\prime\prime}(\bar{\theta})=(v^{\prime}(\theta_{0})+1)f(\bar{\theta})\geq 0$. Denote
-$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+1/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta,$ (B.8)
-Note that $z(\bar{\theta})-F(\bar{\theta})-v^{\prime}(\bar{\theta})=0$, and under IFR,
-$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
-Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}>0$.
-Hence, no exclusion is optimal (i.e., $\theta^{*}_{0}=0$).
+$\begin{aligned} W^{\prime}(\theta_{0})=-({1-F(\theta_{0})})(v^{\prime}(\theta_{0})+1/2+F(\theta_{0})/2)+\frac{f(\theta_{0})}{2}\int_{\theta_{0}}^{\bar{\theta}}({1-F(\theta)})\mathop{}\!\mathrm{d}\theta\end{aligned}$ (B.8)
+The derivative is the same as in the proof of Proposition 4, so the same argument implies that $W^{\prime}(\theta_{0})\leq 0$ for all $\theta_{0}>0$ and hence no exclusion is optimal (i.e., $\theta^{*}_{0}=0$).
 (ii) If $\frac{1-F(\theta)}{f(\theta)}$ is increasing, $s^{*}(\theta)=F(\theta)$ for all $\theta\geq\theta_{0}$, so
@@
 Therefore, no exclusion is optimal because $W^{\prime}(\theta_{0})=-(1-F(\theta_{0}))(F(\theta_{0})+v^{\prime}(\theta))\leq 0$ for all $\theta_{0}>0$. ∎
-B.2 Proof of Section 4
+B.3 Proofs of Section 4
 Proof of Proposition 9.
@@
 ∎
-Because the exclusion level $\theta^{*}_{0}=J^{-1}(0)$ in the standard case of pure intrinsic value coincides with that in the pure status value case in the previous section, the proposition follows directly from the lemma.
+Because the exclusion level $\theta^{*}_{0}=J^{-1}(0)$ in the pure intrinsic value case (Mussa and Rosen, 1978, e.g.,) coincides with that in the pure status value case (i.e., when $v(\theta)=0$) in Proposition 2, Proposition 7 follows directly from the lemma.
 ∎
+Proof of Proposition 8.
+I prove the results for revenue maximization; others can be proven analogously.
+The revenue maximization problem can be written as
+$\max_{{\text{$\tilde{s}$ is feasible}},\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)\tilde{s}(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (B.11)
+Because the objective is linear in $\tilde{s}$, the optimal status allocation $\tilde{s}$ is an extreme point of the feasible set.
+Thus, it is without loss to restrict attention to $\tilde{s}=\phi\circ s$ for some $s\in\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)$
+Consider a relaxed problem where $\tilde{s}(\theta)=\phi(s(\theta))$ for some $s\in\operatorname{\text{MPS}}_{w}(F)$ (instead of $\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)$).
+$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)\phi(s(\theta))\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (B.12)
+(i) Suppose $\phi$ is strictly increasing and convex.
+Then, if $J(\theta)$ is increasing, $J(\theta)\phi(s)$ is convex in $s$ and supermodular in $(s,\theta)$.
+By Fan-Lorentz Theorem, Lemma B.1 still holds (i.e., a finer partition increases the expected revenue).
+Therefore, the solution to the relaxed problem is $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$ for some $\theta_{0}^{*}$.
+Because $s^{*}(\theta)$ is also an extreme point of $\operatorname{\text{MPS}}_{w}(F)$, it is also the solution to the original problem.
+(ii) Suppose $\phi$ is strictly increasing and concave.
+Then, I use the optimal control method to show that $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$ for some $\theta_{0}^{*}$ if $J(\theta)\phi^{\prime}(F(\theta))$ is increasing in $\theta$.
+Define $D(\theta)=\int_{\theta}^{\bar{\theta}}(F(x)-s(x))\mathop{}\!\mathrm{d}F(x)\geq 0$.
+The constraints can be written as $D(\theta)\geq 0$ and $D^{\prime}(\theta)=(s(\theta)-F(\theta))f(\theta)$.
+Thus, the Hamiltonian is
+$H=J(\theta)\phi(s)f(\theta)+\lambda D+\Lambda(s-F)f,$ (B.13)
+where $\lambda(\theta)\geq 0$ is the Lagarangian multiplier on $D(\theta)\geq 0$, and $\Lambda(\theta)$ is the costate variable associated with $D^{\prime}(\theta)$.
+The Pontryagin’s maximum principle implies that
+$\frac{\partial H}{\partial s}=J(\theta)\phi^{\prime}(s(\theta))f(\theta)+\Lambda(\theta)f(\theta)=0,$ (B.14)
+$\frac{\partial H}{\partial D}=\lambda(\theta)=-\dot{\Lambda}(\theta),$ (B.15)
+$\lambda(\theta)D(\theta)=0,\;\lambda(\theta)\geq 0,$ (B.16)
+$\Lambda(\theta_{0})=0,$ (B.17)
+$J(\theta_{0})\phi(s(\theta_{0}))-v^{\prime}(\theta_{0})(1-F(\theta_{0}))+v(\theta_{0})f(\theta_{0})=0.$ (B.18)
+Note that the problem is concave in $s$ because $\phi$ is concave, so the first condition on $s$ is sufficient for optimality.
+Therefore, if $\Lambda(\theta)=J(\theta)\phi^{\prime}(F(\theta))$ is increasing in $\theta$, the solution to the relaxed problem is $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$ for some $\theta_{0}^{*}$.
+Because $s^{*}(\theta)$ is also an extreme point of $\operatorname{\text{MPS}}_{w}(F)$, it is also the solution to the original problem.
+∎
+Proof of Proposition 9.
+Denote by $\widehat{\operatorname{\text{MPS}}}_{w}(F)$ by the set of increasing functions $s\colon\Theta\to[-M,1]$ that satisfies $s\leq\hat{s}$ for some $\hat{s}\colon\Theta\to[0,1]$ such that $\hat{s}\in\operatorname{\text{MPS}}_{w}(F)$.
+$\max_{s\in\widehat{\operatorname{\text{MPS}}}_{w}(F)}\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)(s(\theta)+v^{\prime}(\theta))\right)\mathop{}\!\mathrm{d}F(\theta).$ (B.19)
+Therefore, if $J(\theta)$ is increasing, then $s^{*}=F(\theta)$ for all $\theta\geq J^{-1}(0)$ and $s^{*}(\theta)=-v^{\prime}(\theta)$ for all $\theta<J^{-1}(0)$. ∎